### v2

# TLUE: A Tibetan Language Understanding Evaluation Benchmark

Abstract
Large language models (LLMs) have made tremendous progress in recent years, but low-resource languages, such as Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of LLMs. To address this gap, we present TLUE (A Tibetan Language Understanding Evaluation Benchmark), the first large-scale benchmark for assessing LLMs’ capabilities in Tibetan. TLUE comprises two major components: (1) a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and (2) a safety benchmark covering 7 subdomains. We evaluate a diverse set of state-of-the-art LLMs. Experimental results demonstrate that most LLMs perform below the random baseline, highlighting the considerable challenges LLMs face in processing Tibetan, a low-resource language. TLUE provides an essential foundation for driving future research and progress in Tibetan language understanding and underscores the need for greater inclusivity in LLM development.
TLUE: A Tibetan Language Understanding Evaluation Benchmark
[…]
Figure 1: Overview of the TLUE benchmark.
1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE and SuperGLUE Wang et al. (2018, 2019). However, they largely overlook low-resource languages, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025). Tibetan, spoken by over seven million people, is a key example of such a language, with unique grammatical features and limited annotated data of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023).
Most language understanding benchmarks, including GLUE, SuperGLUE, and BIG-bench, focus on high-resource languages Wang et al. (2018, 2019); Srivastava et al. (2022), leaving low-resource languages like Tibetan without tailored evaluation frameworks. Existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). This paper introduces TLUE: A Tibetan Language Understanding Evaluation Benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation.
Our key contributions are as follows:
  • Developed and released TLUE, the first large-scale Tibetan language evaluation benchmark, covering 67 knowledge subdomains (Ti-MMLU) and 7 safety categories (Ti-SafetyBench).
  • Benchmarked leading LLMs (GPT-4O, Claude-Sonnet, Gemini 1.5, LLaMA 3.1, Qwen2.5, Deepseek) to assess their Tibetan language understanding and safety performance.
  • Conducted qualitative analysis identifying significant limitations in current models, with most performing below the random baseline, emphasizing the need for advancements in low-resource language modeling.
2 Related Work
Language Understanding Benchmarks:
LLMs have been extensively evaluated on multilingual and domain-specific benchmarks, such as XTREME Hu et al. (2020), XGLUE Liang et al. (2020), and MASSIVE FitzGerald et al. (2022). However, these benchmarks predominantly focus on high-resource languages, with little attention given to low-resource languages like Tibetan.
Low-Resource Language Evaluation:
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane $\forall$ et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation. To address this issue, we propose Ti-MMLU, a large-scale Tibetan understanding benchmark inspired by the task design of CMMLU Li et al. (2023), enabling systematic evaluation of model performance in a challenging low-resource setting.
Safety and Ethical Considerations in LLMs:
Recent work has examined the safety and ethical risks of LLMs, particularly in multilingual and low-resource contexts. Studies such as SafetyBench Zhang et al. (2023) and HolisticEval Liang et al. (2023) assess biases, fairness, and robustness in AI systems. However, these benchmarks largely exclude Tibetan and other low-resource languages, leaving significant gaps in understanding how safety concerns manifest in such linguistic environments. Ti-SafetyBench introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE, the first large-scale Tibetan language understanding evaluation benchmark in the world. By offering a comprehensive, multi-domain framework for both language understanding and safety evaluation, TLUE fills a critical gap in the current landscape of LLM benchmarks, enabling systematic assessment of Tibetan language models in a challenging low-resource setting.
3 TLUE
In designing TLUE, we adhere to several key principles to ensure a rigorous and meaningful evaluation of Tibetan LLMs. First, we align our benchmark with existing multilingual evaluation frameworks to facilitate direct comparison with established benchmarks. This allows researchers to assess the performance of Tibetan LLMs relative to models trained in other languages, providing a broader perspective on cross-lingual capabilities Park et al. (2024). Additionally, we ensure that the evaluation data is carefully curated and free from data leakage, maintaining the integrity of benchmark results and preventing unintended memorization effects Doe and Smith (2024).
Figure 2: The curation process of TLUE
Curation process.
In the curation process outlined in Figure 2, the TLUE benchmark follows a structured multi-stage process to ensure the quality and cultural relevance of Tibetan evaluation data. Source datasets, such as CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), are first translated into Tibetan using Claude 3.5 Sonnet Anthropic (2024), which demonstrates strong Tibetan text generation capabilities. The translations undergo rule-based filtering to correct syntactic inconsistencies and structural errors.
Following this, expert human annotators, including two Tibetan language specialists and a team of 15 annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
The final phase includes domain-specific validation, where legal, medical, and scientific content is reviewed by subject-matter experts for accuracy. This multi-layered curation approach ensures TLUE remains a comprehensive and culturally adapted benchmark for Tibetan language model evaluation.
Dataset Sizes.
TLUE consists of 22,963 evaluation questions, divided into Ti-MMLU for knowledge assessment and Ti-SafetyBench for safety evaluation. Ti-MMLU includes 11,528 multiple-choice questions across 67 subjects, covering social sciences, STEM, China-specific topics, humanities, and other domains, enabling a comprehensive evaluation of Tibetan language understanding. Ti-SafetyBench contains 11,435 questions across 7 safety categories, including ethics, bias, health, and privacy, designed to assess LLMs’ ability to handle sensitive content. Dataset sizes and licenses are detailed in Appendix Table 3, adhering to the original dataset licenses where applicable.While Ti-MMLU focuses on broad, multi-domain knowledge evaluation, Ti-SafetyBench targets high-precision safety assessments, together forming a balanced and comprehensive Tibetan language benchmark. The dataset sizes are summarized in Figure 1, which illustrates the comprehensive breakdown of Ti-MMLU and Ti-SafetyBench, highlighting the wide range of subjects and safety categories covered.
Why Extend Existing Benchmarks?
Developing a benchmark for Tibetan, a low-resource language, presents challenges such as limited annotated corpora and linguistic expertise. TLUE extends existing benchmarks via translation, enhanced with human annotation and quality control. Key reasons include:
  • Enabling Cross-Lingual Comparability: A translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
  • Addressing Data Scarcity and Leveraging Established Designs: Due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
  • Ensuring Quality through Controlled Adaptation: TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards.
This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
4 Experiments and Analysis
We use TLUE to evaluate the Tibetan language understanding and safety capabilities of various LLMs. Our experiments cover a diverse range of models, including both open-source and proprietary models.
  • Open-source models. We evaluate several open-source LLMs, including LLaMA-3.1-405B-Instruct Dubey et al. (2024), LLaMA-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-72B-Instruct Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These models represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
  • Proprietary models. We also evaluate several proprietary models, including GPT-4OOpenAI (2024), GPT-3.5-Turbo OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
We investigate several key aspects of LLM performance in Tibetan. We first evaluate all models on Ti-MMLU and Ti-SafetyBench, measuring their Tibetan multi-task understanding and safety capabilities. We then assess the impact of language resource availability by comparing CMMLU (Chinese) and Ti-MMLU(Tibetan) and extend this analysis to safety evaluation by comparing SafetyBench (Chinese) with Ti-SafetyBench. Furthermore, we explore the differences between reasoning-optimized models and chat models, examining whether reasoning enhancements improve low-resource language performance. Finally, we study the effect of model size on Tibetan language understanding, comparing different parameter variants of LLaMA-3.1 and Qwen2.5.
Evaluation Methodology.
We employ zero-shot evaluation to simulate real-world conditions where models must perform tasks without task-specific examples, providing a more accurate assessment of performance in low-resource languages like Tibetan. Appendix Figure 6 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
Evaluation Metrics.
We evaluate model performance on TLUE using two metrics: response rate and accuracy. Response rate measures the proportion of valid responses, while accuracy tracks the proportion of correct answers. For Ti-MMLU, we also compute conditional accuracy, which reflects the model’s performance when giving valid responses (see Appendix). We use two methods for metric calculation:
  • Direct Answer Calculation (DA): Measures if the model selects a valid and correct answer. See Appendix Algorithm 1 for details.
  • Concern All Answer Calculation (CAA): Evaluates the model’s maximum performance by considering all valid options and selecting the correct answer based on remaining choices. See Appendix Algorithm 2 for details.
These methods provide a thorough assessment of the model’s Tibetan language performance, with DA focusing on correct responses and CAA on the model’s potential across all choices.
4.1 Overall Performance on TLUE
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Claude-3.5-Sonnet DA 33.95 28.04 36.32 39.31 34.01 32.09
Gemini-1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7
GPT4O 16.00 12.73 16.36 18.04 16.92 15.96
GPT-3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87
Deepseek-v3 29.51 23.57 31.97 33.65 29.92 28.44
Llama3.1-405B-Instruct 25.08 23.88 24.25 25.58 27.62 24.07
Llama3.1-70B-Instruct 23.73 23.16 23.2 26.2 24.65 21.45
Qwen2.5-32B-Instruct 13.94 12.63 14.98 14.92 13.71 13.44
Qwen2.5-72B-Instruct 7.27 6.07 7.98 7.52 7.74 7.02
Qwen2.5-7B-Instruct 1.8 2.94 1.87 1.63 0.9 1.68
Llama3.1-8B-Instruct 5.47 5.48 5.56 5.99 5.46 4.86
Deepseek-r1 15.74 13.87 14.24 18.06 13.94 18.58
O1-mini 6.14 6.15 6.33 7.17 6.06 4.98
Claude-3.5-Sonnet CAA 35.63 30.88 37.47 40.58 35.26 33.96
Gemini-1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81
GPT4O 17.51 14.25 17.71 19.69 18.46 17.45
GPT-3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16
Deepseek-v3 32.16 27.03 34.58 36.26 32.00 30.94
Llama3.1-405B-Instruct 25.28 24.10 24.50 25.87 27.73 24.22
Llama3.1-70B-Instruct 23.79 23.22 23.24 26.31 24.65 21.52
Qwen2.5-32B-Instruct 18.56 16.66 20.3 19.72 17.47 18.67
Qwen2.5-72B-Instruct 16.50 15.73 17.88 17.00 15.84 16.04
Qwen2.5-7B-Instruct 14.59 13.92 13.66 16.34 14.57 14.46
Llama3.1-8B-Instruct 7.44 7.95 7.54 7.38 7.41 6.92
Deepseek-r1 27.45 21.01 25.99 32.72 26.44 31.08
O1-mini 9.67 9.69 9.80 10.14 9.68 9.02
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 1: Accuracy results of different LLMs on the Ti-MMLU dataset using two evaluation methods: Concern All Answer (CAA) and Direct Answer (DA).
Tables 1 and 2 show results on Ti-MMLU and Ti-SafetyBench, evaluating Tibetan language understanding and safety.The detial data is released in Appendix Table 14, 15, 16, 17
Performance on Ti-MMLU.
Most models perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary models, the GPT series underperforms, whereas open-source models like Deepseek-v3 exceed the baseline. For reasoning-optimized models, Deepseek-r1 surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
The performance gap between proprietary and open-source models is minimal, with Deepseek-v3 and Gemini-1.5-Flash performing comparably to Claude-3.5-Sonnet. Notably, GPT-3.5-Turbo underperforms relative to Llama3.1-8B-Instruct, suggesting that pretraining data, adaptation, and optimization significantly impact Tibetan language performance.
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, Llama3.1-405B-Instruct excels in Other, and Qwen2.5-72b-Instruct in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
Performance on Ti-SafetyBench.
Most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet achieves the highest accuracy (58.5% in CAA), whereas GPT-4O and GPT-3.5-Turbo underperform, with the latter significantly below random.
Among open-source models, Deepseek-v3 surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. Llama3.1-405B-Instruct approaches the baseline, while smaller models like Llama3.1-8B-Instruct and Qwen2.5-72B-Instruct perform poorly. Reasoning-optimized models such as Deepseek-r1 excel in CAA but struggle in DA, whereas O1-mini consistently underperforms.
These findings suggest that safety alignment in Tibetan remains challenging, with pretraining data, adaptation strategies, and model architecture playing a more critical role than accessibility.
Model Eval Method Avg. OFF UB PH MH IA EM PP
Claude-3.5-Sonnet DA 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini-1.5-Flash 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
GPT-4O 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT-3.5-turbo 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
Deepseek-v3 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
Llama3.1-405B-Instruct 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
Llama3.1-70B-Instruct 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
Qwen2.5-72B-Instruct 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen2.5-32B-Instruct 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen2.5-7B-Instruct 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
Llama3.1-8B-Instruct 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
Deepseek-r1 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
O1-mini 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Claude-3.5-Sonnet CAA 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini-1.5-Flash 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
GPT-4O 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT-3.5-turbo 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
Deepseek-v3 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
Llama3.1-405B-Instruct 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
Llama3.1-70B-Instruct 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
Qwen2.5-72B-Instruct 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen2.5-32B-Instruct 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen2.5-7B-Instruct 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Llama3.1-8B-Instruct 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
Deepseek-r1 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
O1-mini 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 2: Accuracy results of different LLMs on the Ti-SafetyBench dataset using two evaluation methods: Concern All Answer (CAA) and Direct Answer (DA). The full form of the 7 categories is provided in Appendix Table 4.
4.2 High-Resource vs. Low-Resource Performance
To examine the performance disparity between high-resource and low-resource languages, we compare model accuracy on CMMLU (Chinese) Li et al. (2023) and Ti-MMLU (Tibetan) using the Concern All Answer (CAA) evaluation. Figure 3(a) illustrates performance differences across domains, while Appendix Table 5 presents detailed accuracy comparisons.
Performance Degradation.
All models experience substantial accuracy drops from CMMLU to Ti-MMLU. Qwen2.5-72B-Instruct declines from 84.70% $\rightarrow$ 16.50%, GPT-4 from 68.90% $\rightarrow$ 17.51%, and ChatGPT from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%). This underscores the difficulty of adapting LLMs to low-resource languages.
Domain Performance Shifts.
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU, models excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. Notable shifts include:
  • Qwen2.5-72B-Instruct achieves the highest accuracy in "Other" (87.35%) on CMMLU but shifts to "Humanities" (17.88%) on Ti-MMLU.
  • Llama3.1-70B-Instruct transitions from "Other" (74.72%) in CMMLU to "Social Sciences" (26.31%) in Ti-MMLU.
  • GPT-4’s best-performing category changes from "Other" (73.16%) in CMMLU to "Social Sciences" (19.69%) in Ti-MMLU(GPT-4O).
These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
STEM as the Weakest Domain.
STEM remains the most challenging domain in both benchmarks. It consistently ranks lowest in CMMLU and Ti-MMLU, except for Llama3.1-70B-Instruct, which performs worst in "China-specific" tasks. This indicates significant difficulties in mathematical and technical reasoning in Tibetan.
Ranking Inconsistencies.
Model rankings in CMMLU do not consistently translate to Ti-MMLU. For instance, Llama3.1-70B-Instruct ranks below Qwen2.5-72B-Instruct in CMMLU but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
Implications for Low-Resource Adaptation.
The substantial performance gap underscores the need for improved Tibetan data coverage in pretraining and fine-tuning. While LLMs excel in high-resource languages, their struggles in Tibetan highlight the critical role of data availability and adaptation strategies in enhancing low-resource language understanding.
4.3 Safety in Low-Resource Languages
(a) Performance degradation from CMMLU (Chinese) to Ti-MMLU (Tibetan).
(b) Performance degradation from SafetyBench (Chinese) to Ti-SafetyBench (Tibetan)
Figure 3: Comparison of model performance on Chinese Benchmarks and TLUE across different tasks.
To evaluate model safety in low-resource languages, we compare performance on SafetyBench (Chinese) Zhang et al. (2023) and Ti-SafetyBench (Tibetan) using the Concern All Answer (CAA) evaluation. Figure 3(b) illustrates accuracy drops across safety categories, while Appendix Table 6 provides detailed performance breakdowns. These results highlight the challenges of maintaining safety alignment in Tibetan.
Performance Degradation in Tibetan.
Models show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 achieves 89.2% on SafetyBench, but GPT-4O, despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
Variability in Safety Categories.
Performance degradation varies across safety categories. GPT-4 excels in Physical Health (PH) and Ethics and Morality (EM) on SafetyBench, whereas GPT-4O, despite being more advanced, performs best in Illegal Activities (IA) on Ti-SafetyBench, highlighting the challenges of transferring safety alignment to Tibetan.
Similarly, Qwen-chat-7B outperforms random in EM and IA on SafetyBench, but Qwen2.5-7B-Instruct retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
GPT-3.5-Turbo falls well below the random baseline on Ti-SafetyBench, while GPT-4O and Qwen2.5-7B-Instruct exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
Challenges in Low-Resource Safety Alignment.
The sharp decline in performance on Ti-SafetyBench underscores the challenge of maintaining safety alignment in Tibetan. While models perform well on SafetyBench (Chinese), most fall below or barely match the baseline on Ti-SafetyBench. Even top-performing models struggle with safety generalization, highlighting the need for targeted fine-tuning and enhanced multilingual adaptation in low-resource environments.
4.4 Reasoning vs. Chat Models
We compare reasoning-optimized and chat models on TLUE, analyzing response behavior, accuracy, and conditional accuracy. The results are summarized in Appendix Table 7811, with a broader comparison provided on Ti-MMLU (Appendix Table 9) and Ti-SafetyBench (Appendix Table 10), covering overall accuracy across knowledge and safety domains.
Response Rate and Output Behavior.
Reasoning models, compared to chat models, tend to evaluate all answer choices before selecting the correct one, resulting in higher response rates for CAA than DA (Figure4(a), 4(b),Appendix Table 78). This behavior stems from a step-by-step evaluation process. When evaluating reasoning models, intermediate reasoning steps are filtered, and only the final output is considered, as outlined in Appendix Algorithm 3 and 4.
Impact of Response Rate on Performance.
Lower response rates notably affect reasoning models’ performance, highlighting the inherent difficulty in responding to Tibetan prompts. The gap between DA and CAA accuracy underscores this challenge, with DA tasks requiring direct answer generation (Figure4(a), 4(b), Appendix Table 78). Reasoning models, unlike chat models, struggle more with Tibetan prompts due to their emphasis on logical inference over language generation.
Generalization of Reasoning Models in Low-Resource Languages.
As shown in Appendix Table 11, Deepseek-r1 consistently outperforms Deepseek-v3 in conditional accuracy. Additionally, O1-mini outperforms GPT-4O in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini underperforms compared to GPT-4O in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
(a) Response rate, accuracy, and conditional accuracy on Ti-MMLU.
(b) Response rate and accuracy on Ti-SafetyBench.
Figure 4: Response rate, accuracy, and conditional accuracy for reasoning-optimized and chat models on TLUE.
4.5 Model Scale and Low-Resource Performance
We investigate the effect of model scale on Tibetan language understanding and safety alignment using Ti-MMLU (Figure 5(a) , Appendix Table 12) and Ti-SafetyBench (Figure 5(b), Appendix Table 13).
Effect of Scale on Tibetan Understanding.
Larger models generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LLaMA-3.1-405B-Instruct slightly outperforms LLaMA-3.1-70B-Instruct, while Qwen2.5-32B-Instruct surpasses Qwen2.5-72B-Instruct. Notably, Qwen2.5-7B-Instruct achieves comparable or superior accuracy to Qwen2.5-72B-Instruct in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
Impact of Scale on Safety Alignment.
Larger models show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen2.5-72B-Instruct significantly outperforms Qwen2.5-32B-Instruct, indicating that scaling benefits safety alignment more than general understanding. In contrast, LLaMA-3.1-70B-Instruct shows only marginal improvement over LLaMA-3.1-8B-Instruct, suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
Challenges in Scaling for Low-Resource Languages.
While larger models generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller models, such as Qwen2.5-7B-Instruct, achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.
(a) Average accuracy on Ti-MMLU
(b) Average accuracy on Ti-SafetyBench
Figure 5: Average accuracy on TLUE across different model scales for LLaMA-3.1 and Qwen2.5.
5 Conclusion
This paper presents TLUE, a benchmark for evaluating large language models (LLMs) in Tibetan, covering general knowledge understanding (Ti-MMLU) and safety alignment (Ti-SafetyBench). Our findings show that both proprietary and open-source models struggle in Tibetan, performing below or near the random baseline. We identify key challenges, including significant performance drops when transitioning from high-resource languages to Tibetan, limitations in current multilingual training, and the need for enhanced Tibetan language resources. While open-source models show promising domain-specific improvements, reasoning-optimized models demonstrate better generalization but face difficulties in following Tibetan prompts. Model scaling provides inconsistent benefits, with architecture, data quality, and fine-tuning strategies proving more crucial than model size. These results underscore the need for improved pretraining and targeted fine-tuning, with TLUE serving as a foundation for future low-resource language modeling research, promoting inclusivity and robustness in LLM development.
6 Limitations
The limitation of TLUE is the limited coverage of Tibetan cultural and folk knowledge. While the benchmark includes diverse linguistic and domain-specific tasks, it lacks evaluation datasets tailored to Tibetan folklore, traditional customs, and region-specific knowledge. As cultural understanding is crucial for real-world applications, future work should expand TLUE to better assess LLMs’ capabilities in these areas.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.
Appendix A Appendix
Dataset # Samples License
Ti-MMLU 11.5K CC-BY-NC-SA 4.0
Ti-SafetyBench 11.4K Apache License 2.0
Table 3: Number of samples and license information for each of the datasets in the TLUE benchmark.
Figure 6: Prompt and Examples of TLUE
Abbreviation Full - form
OFF Offensiveness
UB Unfairness and Bias
PH Physical Health
MH Mental Health
IA Illegal Activities
EM Ethics and Morality
PP Privacy and Property
Table 4: Abbreviations of different types of tasks in Ti-SafetyBench.
Algorithm 1 Algorithm for Direct Answer Calculation (DA) with Variable Choice Count
1:procedure ExtractDirectAnswer(model_result, exclude_list, option_num)
2:     $valid\_choices\_map\leftarrow\{\text{`2''}:\{A,B\},\text{3''}:\{A,B,C\},\text{4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
9:     end for
10:     $found\_choices\leftarrow[choice\in valid\_choices\mid choice\in model\_result]$
11:     if $|found\_choices|=1$ then
12:         return $found\_choices[0]$triangleright Return extracted answer if unique
13:     else
14:         return $\emptyset$ $\triangleright$ Return empty if no valid answer is found
15:     end if
16:end procedure
Algorithm 2 Algorithm for Concern All Answer Calculation (CAA) with Variable Choice Count
1:procedure ExtractCAAAnswer(model_result, exclude_list, option_num)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
9:     end for
10:     Step 1: Extract all valid choices in order
11:     $all\_choices\leftarrow[ch\in model\_result\mid ch\in valid\_choices]$
12:     Step 2: Remove full valid choice combinations (unordered)
13:     $filtered\_choices\leftarrow[]$, $buffer\leftarrow[]$
14:     for each $ch$ in $all\_choices$ do
15:         Append $ch$ to $buffer$
16:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
17:              Clear $buffer$triangleright Remove one full valid choice combination
18:         else if $|buffer|>option\_num$ then
19:              Append first elements of $buffer$ to $filtered\_choices$
20:              Keep only last $(option\_num-1)$ elements in $buffer$
21:         end if
22:     end for
23:     Append remaining elements in $buffer$ to $filtered\_choices$
24:     Step 3: Determine final answer
25:     $unique\_choices\leftarrow set(filtered\_choices)$
26:     if $|unique\_choices|=1$ then
27:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains
28:     else
29:         return $\emptyset$ $\triangleright$ Return empty if multiple choices remain
30:     end if
31:end procedure
Algorithm 3 Algorithm for Direct Answer Extraction with Reasoning Filter
1:procedure ExtractDirectAnswer(model_result, exclude_list, option_num, model_name)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Filter model-specific reasoning content
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “Deepseek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
11:     end if
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
14:     end for
15:     $found\_choices\leftarrow[choice\in valid\_choices\mid choice\in model\_result]$
16:     if $|found\_choices|=1$ then
17:         return $found\_choices[0]$triangleright Return extracted answer if unique
18:     else
19:         return $\emptyset$ $\triangleright$ Return empty if no valid answer is found
20:     end if
21:end procedure
Algorithm 4 Algorithm for Concern All Answer Calculation (CAA) with Reasoning Filter
1:procedure ExtractCAAAnswer(model_result, exclude_list, option_num, model_name)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Apply model-specific reasoning content filtering
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “Deepseek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
11:     end if
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
14:     end for
15:     Step 1: Extract all valid choices in order
16:     $all\_choices\leftarrow[ch\in model\_result\mid ch\in valid\_choices]$
17:     Step 2: Remove full valid choice combinations (unordered)
18:     $filtered\_choices\leftarrow[]$, $buffer\leftarrow[]$
19:     for each $ch$ in $all\_choices$ do
20:         Append $ch$ to $buffer$
21:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
22:              Clear $buffer$triangleright Remove one full valid choice combination
23:         else if $|buffer|>option\_num$ then
24:              Append first elements of $buffer$ to $filtered\_choices$
25:              Keep only last $(option\_num-1)$ elements in $buffer$
26:         end if
27:     end for
28:     Append remaining elements in $buffer$ to $filtered\_choices$
29:     Step 3: Determine final answer
30:     $unique\_choices\leftarrow set(filtered\_choices)$
31:     if $|unique\_choices|=1$ then
32:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains
33:     else
34:         return $\emptyset$ $\triangleright$ Return empty if multiple choices remain
35:     end if
36:end procedure
Model Benchmark Avg. STEM Humanities Social Sciences Other China specific
Qwen2.5-72B-Instruct CMMLU 84.70 80.67 87.00 84.66 87.35 83.21
Ti - CMMLU 16.50 15.73 17.88 17.00 15.84 16.04
Llama3.1-70B-Instruct CMMLU 69.01 61.60 71.44 69.42 74.72 63.79
Ti - CMMLU 23.79 23.22 23.24 26.31 24.65 21.52
GPT4 CMMLU 68.90 63.16 69.19 70.26 73.16 63.47
GPT4O Ti - CMMLU 17.51 14.25 17.71 19.69 18.46 17.45
ChatGPT CMMLU 53.22 44.80 53.61 54.22 59.95 49.74
GPT - 3.5 - Turbo Ti - CMMLU 3.42 3.82 3.35 3.68 3.09 3.16
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 5: Comparison of model performance on CMMLU (Chinese) and Ti-MMLU(Tibetan). The table shows a significant performance drop in Ti-MMLU, with most models falling below the random baseline (25%).
Model Eval Method Avg. OFF UB PH MH IA EM PP
GPT4 SafetyBench 89.2 85.4 76.4 95.5 94.1 92.5 92.6 92.5
GPT4O Ti - SafetyBench 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT-3.5-Turbo SafetyBench 80.4 76.1 68.7 78.4 89.7 87.3 78.5 87.9
Ti - SafetyBench 11.6 12.5 16 11.4 10.3 8.2 10.8 11.8
Qwen-chat-7B SafetyBench 77.4 72.4 64.4 71.5 89.3 84.9 78.2 82.4
Qwen2.5-7B-Instruct Ti - SafetyBench 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 6: Performance comparison of models on SafetyBench (Chinese) and Ti-SafetyBench (Tibetan) across various safety categories. Models show a substantial drop in accuracy on Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
Model Eval Method Response Rate Accuracy Conditional Accuracy
Deepseek-v3 DA 76.99% 29.51 38.3
CAA 84.73% 32.16 37.9
Deepseek-r1 DA 34.84% 15.74 43.75
CAA 57.45% 27.45 47.01
GPT4O DA 51.68% 16 30.93
CAA 57.47% 17.51 30.44
O1mini DA 22.71% 6.14 27.4
CAA 35% 9.67 27.74
Table 7: Response rate, accuracy, and conditional accuracy for reasoning-optimized and chat models on Ti-MMLU.
Model Eval Method Response Rate Accuracy
Deepseek-v3 DA 61.34% 37.4
CAA 78.51% 48.3
Deepseek-r1 DA 40.80% 24.3
CAA 71.86% 46.8
GPT4O DA 61.41% 31.1
CAA 65.45% 32.9
O1mini DA 28.48% 10.9
CAA 38.31% 15.1
Table 8: Response rate and accuracy for reasoning-optimized and chat models on Ti-SafetyBench.
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Deepseek-v3 DA 29.51 23.57 31.97 33.65 29.92 28.44
CAA 32.16 27.03 34.58 36.26 32.00 30.94
Deepseek-r1 DA 15.74 13.87 14.24 18.06 13.94 18.58
CAA 27.45 21.01 25.99 32.72 26.44 31.08
GPT4O DA 16.00 12.73 16.36 18.04 16.92 15.96
CAA 17.51 14.25 17.71 19.69 18.46 17.45
O1-mini DA 6.14 6.15 6.33 7.17 6.06 4.98
CAA 9.67 9.69 9.80 10.14 9.68 9.02
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 9: Accuracy comparison of reasoning-optimized and chat models on Ti-MMLUacross different knowledge domains.
Model Eval Method Avg. OFF UB PH MH IA EM PP
Deepseek-v3 DA 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
CAA 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
Deepseek-r1 DA 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
CAA 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
GPT-4O DA 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
CAA 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
O1-mini DA 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
CAA 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 10: Accuracy comparison of reasoning-optimized and chat models on Ti-SafetyBench across different safety categories.
Model Eval Method Conditional Accuracy (Avg.) STEM Humanities Social Sciences Other China specific
Deepseek-v3 DA 38.3 38.52 39.7 39.61 38.2 35.46
CAA 37.9 37.93 38.89 39.45 38.2 35.03
Deepseek-r1 DA 43.75 44.81 38.57 47.39 44.24 43.73
CAA 47.01 46.97 44.68 49.79 48.23 45.37
GPT4O DA 30.93 29.07 30.62 32.59 31.93 30.44
CAA 30.44 28.86 30.04 31.85 31.51 29.96
O1mini DA 27.4 33.24 27.07 26.79 26.1 23.78
CAA 27.74 31.48 28.53 26.01 26.26 26.43
Table 11: Conditional accuracy across different knowledge domains for reasoning-optimized and chat models on Ti-MMLU.
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Llama3.1-405B-Instruct DA 25.08 23.88 24.25 25.58 27.62 24.07
CAA 25.28 24.10 24.50 25.87 27.73 24.22
Llama3.1-70B-Instruct DA 23.73 23.16 23.2 26.2 24.65 21.45
CAA 23.79 23.22 23.24 26.31 24.65 21.52
Llama3.1-8B-Instruct DA 5.47 5.48 5.56 5.99 5.46 4.86
CAA 7.44 7.95 7.54 7.38 7.41 6.92
Qwen2.5-32B-Instruct DA 13.94 12.63 14.98 14.92 13.71 13.44
CAA 18.56 16.66 20.3 19.72 17.47 18.67
Qwen2.5-72B-Instruct DA 7.27 6.07 7.98 7.52 7.74 7.02
CAA 16.50 15.73 17.88 17.00 15.84 16.04
Qwen2.5-7B-Instruct DA 1.8 2.94 1.87 1.63 0.9 1.68
CAA 14.59 13.92 13.66 16.34 14.57 14.46
Random - 25.00 25.00 25.00 25.00 25.00 25.00
Table 12: Accuracy comparison of different model scales on Ti-MMLU across various knowledge domains.
Model Eval Method Avg. OFF UB PH MH IA EM PP
Llama3.1-405B-Instruct DA 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
CAA 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
Llama3.1-70B-Instruct DA 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
CAA 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
Llama3.1-8B-Instruct DA 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
Qwen2.5-72B-Instruct DA 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
CAA 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen2.5-32B-Instruct DA 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
CAA 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen2.5-7B-Instruct DA 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
CAA 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 13: Accuracy comparison of different model scales on Ti-SafetyBench across various knowledge domains.
Task Name GPT-3.5-turbo GPT4O Claude-3.5-Sonnet Gemini-1.5-flash Deepseek-v3 Llama3.1-8B-Instruct Llama3.1-70B-Instruct
agronomy 2.96 15.38 31.95 30.18 33.14 5.33 15.98
anatomy 1.35 13.51 27.03 27.7 23.65 4.05 29.05
ancient_chinese 3.66 13.41 30.49 26.22 26.22 4.88 14.02
arts 5.0 16.25 33.75 36.25 35.62 7.5 21.25
astronomy 3.03 11.52 35.76 23.03 30.91 7.27 20.0
business_ethics 3.35 22.01 42.11 38.76 33.97 6.22 23.44
chinese_civil_service_exam 3.12 13.12 24.38 21.25 28.75 5.62 19.38
chinese_driving_rule 2.29 19.08 50.38 49.62 39.69 4.58 33.59
chinese_food_culture 2.94 19.12 27.94 34.56 29.41 4.41 29.41
chinese_foreign_policy 0.93 20.56 47.66 38.32 34.58 12.15 32.71
chinese_history 2.17 13.62 32.2 23.22 37.46 5.26 26.32
chinese_literature 2.45 18.63 23.04 28.43 29.41 7.35 19.61
chinese_teacher_qualification 2.23 17.32 45.81 36.87 25.14 6.15 22.91
clinical_knowledge 4.64 16.03 24.89 18.99 9.7 6.33 18.99
college_actuarial_science 4.72 6.6 18.87 17.92 25.47 8.49 17.92
college_education 6.54 18.69 42.99 42.06 39.25 11.21 27.1
college_engineering_hydrology 5.66 21.7 36.79 33.02 31.13 6.6 29.25
college_law 3.7 18.52 26.85 18.52 19.44 10.19 22.22
college_mathematics 6.67 11.43 20.95 19.05 28.57 8.57 18.1
college_medical_statistics 5.66 14.15 41.51 34.91 28.3 11.32 24.53
college_medicine 1.47 16.12 26.37 21.25 23.81 8.06 22.71
computer_science 4.41 14.71 25.49 29.41 29.41 8.33 24.51
computer_security 3.51 22.81 44.44 32.16 30.99 8.77 30.99
conceptual_physics 0.68 16.33 41.5 27.89 25.85 13.61 23.13
construction_project_management 2.16 14.39 38.13 29.5 28.06 8.63 12.95
economics 1.89 13.21 38.99 33.33 39.62 11.32 32.7
education 3.07 18.4 42.33 37.42 36.81 9.82 19.63
electrical_engineering 1.74 19.19 31.98 29.65 27.91 8.14 29.65
elementary_chinese 3.17 12.3 25.79 23.41 32.54 6.35 13.89
elementary_commonsense 6.06 16.16 39.9 32.83 35.86 6.06 24.24
elementary_information_and_technology 5.46 24.37 42.02 41.6 37.82 10.08 29.41
elementary_mathematics 7.39 13.04 36.09 34.78 26.09 6.96 17.39
ethnology 3.7 25.93 34.07 37.78 34.81 5.93 23.7
food_science 2.1 20.28 34.27 25.17 37.76 5.59 28.67
genetics 1.7 15.91 30.68 28.41 25.57 7.39 23.86
global_facts 2.01 19.46 34.9 31.54 38.26 8.05 19.46
high_school_biology 5.33 17.16 20.12 24.26 22.49 9.47 18.34
high_school_chemistry 4.55 12.12 20.45 12.88 16.67 13.64 19.7
high_school_geography 0.85 25.42 40.68 27.97 31.36 7.63 23.73
high_school_mathematics 6.71 12.2 34.15 25.0 37.2 7.32 20.12
high_school_physics 0.91 18.18 39.09 31.82 32.73 4.55 30.0
high_school_politics 4.2 20.28 30.07 23.08 30.77 6.29 21.68
human_sexuality 3.17 19.05 35.71 34.92 43.65 6.35 30.95
international_law 3.24 19.46 36.22 34.59 25.41 6.49 21.08
journalism 3.49 19.19 39.53 36.63 33.14 5.81 26.74
jurisprudence 2.92 19.71 39.17 30.41 37.71 9.25 21.65
legal_and_moral_basis 3.27 25.7 60.75 50.93 53.27 7.94 35.51
logical 3.25 15.45 38.21 28.46 40.65 5.69 26.02
machine_learning 0.82 11.48 29.51 27.87 22.13 6.56 21.31
management 4.76 20.0 40.48 39.05 33.81 6.19 24.76
marketing 4.44 20.56 45.0 43.33 39.44 7.22 31.11
marxist_theory 1.06 21.16 48.68 38.62 41.27 10.05 32.28
modern_chinese 5.17 18.97 24.14 22.41 20.69 8.62 12.93
nutrition 2.07 11.72 37.24 37.24 31.72 6.9 20.0
philosophy 3.81 19.05 44.76 38.1 42.86 7.62 28.57
professional_accounting 5.14 20.0 41.14 37.71 33.71 6.86 20.57
professional_law 1.9 17.06 27.96 19.43 29.38 5.69 17.06
professional_medicine 3.46 13.83 23.67 18.88 24.47 10.9 17.82
professional_psychology 4.31 21.98 37.07 34.05 37.93 6.03 25.0
public_relations 2.87 11.49 32.18 33.33 36.78 5.17 24.14
security_study 5.19 25.19 42.96 37.78 38.52 7.41 30.37
sociology 1.33 18.58 43.81 36.28 37.61 7.52 26.11
sports_science 4.24 23.64 33.33 33.33 31.52 7.88 24.85
traditional_chinese_medicine 3.24 17.84 27.03 28.11 24.86 8.65 17.3
virology 3.55 13.02 34.91 26.04 25.44 2.96 27.81
world_history 4.97 15.53 36.02 28.57 30.43 7.45 24.84
world_religions 5.0 13.12 45.62 31.87 39.38 5.0 21.25
Average 3.47 17.33 35.25 30.87 31.76 7.49 23.68
Table 14: Accuracy on 67 subjects of Ti-MMLU(CAA).The Category Mapping is in 18
Task Name Llama3.1-405B-Instruct Qwen2.5-7B-Instruct Qwen2.5-32B-Instruct Qwen2.5-72B-Instruct O1-mini Deepseek-r1
agronomy 26.04 12.43 18.34 10.65 8.28 28.99
anatomy 20.95 16.89 10.81 11.49 5.41 15.54
ancient_chinese 15.24 17.68 21.34 11.59 10.98 25.61
arts 23.75 13.12 26.25 13.12 6.88 29.38
astronomy 26.06 13.94 16.36 11.52 16.36 24.24
business_ethics 24.4 15.79 14.83 10.05 10.53 30.62
chinese_civil_service_exam 20.62 11.25 14.37 11.88 5.62 25.0
chinese_driving_rule 25.19 16.03 16.79 12.98 12.21 44.27
chinese_food_culture 24.26 11.76 17.65 15.44 6.62 36.76
chinese_foreign_policy 29.91 17.76 18.69 14.02 9.35 28.04
chinese_history 25.39 12.69 17.96 13.93 6.81 27.55
chinese_literature 21.57 15.2 19.61 14.71 13.73 26.47
chinese_teacher_qualification 34.08 15.64 20.11 13.41 10.06 41.34
clinical_knowledge 24.89 10.97 12.24 6.75 7.59 17.3
college_actuarial_science 16.04 5.66 17.92 12.26 8.49 24.53
college_education 39.25 14.02 24.3 15.89 14.95 32.71
college_engineering_hydrology 28.3 15.09 20.75 8.49 12.26 24.53
college_law 22.22 11.11 17.59 12.04 9.26 32.41
college_mathematics 18.1 13.33 10.48 13.33 3.81 9.52
college_medical_statistics 27.36 16.98 25.47 9.43 15.09 5.66
college_medicine 19.05 10.26 12.09 9.16 6.59 12.45
computer_science 28.92 10.78 15.69 9.8 11.27 20.1
computer_security 30.41 16.37 17.54 9.94 12.87 25.15
conceptual_physics 27.89 14.29 15.65 13.61 10.88 30.61
construction_project_management 23.74 16.55 19.42 12.23 7.91 28.06
economics 20.13 15.72 19.5 6.29 6.29 39.62
education 28.83 14.11 15.34 14.11 7.36 31.29
electrical_engineering 26.16 14.53 12.79 14.53 8.14 29.65
elementary_chinese 17.06 9.92 19.44 15.87 7.54 29.76
elementary_commonsense 28.28 14.14 20.2 9.6 8.08 36.36
elementary_information_and_technology 33.19 13.45 18.07 11.34 10.5 23.53
elementary_mathematics 28.7 15.22 22.61 14.78 6.52 44.35
ethnology 25.93 15.56 22.22 13.33 12.59 41.48
food_science 32.87 14.69 18.88 11.19 6.29 25.17
genetics 21.59 11.93 13.64 6.25 10.8 13.64
global_facts 20.81 10.74 18.79 9.4 10.07 18.79
high_school_biology 20.71 8.88 12.43 8.28 7.69 8.28
high_school_chemistry 19.7 7.58 12.12 6.06 9.09 6.06
high_school_geography 22.88 16.95 14.41 11.86 7.63 32.2
high_school_mathematics 24.39 26.83 24.39 18.9 9.15 27.44
high_school_physics 23.64 20.0 20.91 10.91 9.09 33.64
high_school_politics 26.57 15.38 18.18 10.49 6.99 26.57
human_sexuality 27.78 15.87 19.05 15.08 10.32 32.54
international_law 22.16 14.05 25.41 16.76 10.27 5.95
journalism 26.74 22.09 16.28 12.21 7.56 30.81
jurisprudence 27.25 15.09 19.22 11.44 9.49 32.36
legal_and_moral_basis 35.98 24.3 21.96 14.02 8.88 53.27
logical 30.08 12.2 23.58 12.2 10.57 26.83
machine_learning 24.59 9.84 12.3 6.56 6.56 18.03
management 24.76 12.86 20.0 7.14 14.76 39.52
marketing 32.22 13.89 24.44 17.22 12.22 37.22
marxist_theory 23.81 16.4 14.81 16.93 10.05 40.21
modern_chinese 19.83 14.66 20.69 10.34 6.03 29.31
nutrition 27.59 17.24 12.41 7.59 13.79 20.69
philosophy 27.62 14.29 20.0 10.48 13.33 1.9
professional_accounting 22.86 18.86 19.43 12.0 9.14 30.86
professional_law 19.43 10.9 16.59 10.43 10.43 25.12
professional_medicine 18.09 11.44 16.76 8.24 10.64 19.15
professional_psychology 21.12 11.21 23.28 13.79 9.05 33.19
public_relations 24.14 17.82 20.69 13.79 10.34 29.89
security_study 26.67 20.0 20.0 14.07 10.37 25.93
sociology 25.22 16.81 19.47 11.06 8.85 30.09
sports_science 34.55 12.12 25.45 13.33 11.52 29.7
traditional_chinese_medicine 21.08 12.97 16.76 13.51 12.43 19.46
virology 26.63 14.79 18.93 9.47 14.2 21.3
world_history 24.22 16.15 22.36 16.15 11.8 32.3
world_religions 28.12 16.25 18.75 11.25 5.62 40.62
Average 25.19 14.53 18.43 11.88 9.64 27.18
Table 15: Accuracy on 67 subjects of Ti-MMLU(CAA) continue
Task Name GPT-3.5-turbo GPT4O Claude-3.5-Sonnet Gemini-1.5-flash Deepseek-v3 Llama3.1-8B-Instruct Llama3.1-70B-Instruct
agronomy 0.59 13.02 31.36 30.18 30.18 4.73 15.98
anatomy 1.35 12.16 26.35 27.7 19.59 2.7 29.05
ancient_chinese 1.83 12.2 28.05 26.22 22.56 3.05 13.41
arts 4.38 15.0 33.12 36.25 30.0 5.0 21.25
astronomy 2.42 10.3 34.55 21.82 26.06 6.67 20.0
business_ethics 1.91 20.57 40.67 36.84 30.14 5.74 23.44
chinese_civil_service_exam 0.62 10.0 19.38 18.12 25.0 3.75 19.38
chinese_driving_rule 0.76 16.79 50.38 49.62 38.93 3.82 33.59
chinese_food_culture 2.21 17.65 27.21 33.82 27.94 4.41 29.41
chinese_foreign_policy 0.0 20.56 42.99 36.45 32.71 10.28 32.71
chinese_history 1.55 11.46 30.65 21.98 31.58 4.64 26.32
chinese_literature 1.47 17.65 22.55 28.43 27.45 4.9 19.61
chinese_teacher_qualification 2.23 16.76 44.13 35.75 24.02 4.47 22.91
clinical_knowledge 3.38 15.61 24.89 17.72 9.28 4.22 18.99
college_actuarial_science 2.83 4.72 16.98 16.98 22.64 7.55 17.92
college_education 3.74 18.69 42.06 41.12 36.45 8.41 27.1
college_engineering_hydrology 3.77 19.81 36.79 33.02 30.19 2.83 29.25
college_law 1.85 15.74 25.93 18.52 16.67 6.48 22.22
college_mathematics 2.86 9.52 17.14 16.19 20.0 5.71 18.1
college_medical_statistics 2.83 14.15 41.51 33.02 22.64 3.77 24.53
college_medicine 0.73 15.38 26.01 21.25 22.71 5.49 22.71
computer_science 2.45 12.75 25.0 28.43 27.45 4.9 24.51
computer_security 1.75 20.47 39.18 30.41 29.24 7.6 30.99
conceptual_physics 0.68 14.97 35.37 25.85 22.45 10.2 23.13
construction_project_management 1.44 12.95 35.97 28.06 25.9 5.04 12.95
economics 1.26 11.95 37.11 32.7 37.74 8.18 32.7
education 2.45 15.95 41.1 36.2 34.36 6.75 19.63
electrical_engineering 1.74 18.6 30.81 29.65 26.16 5.81 29.65
elementary_chinese 2.38 11.9 24.21 23.41 29.37 4.76 13.49
elementary_commonsense 3.54 15.66 39.9 32.83 32.83 3.54 24.24
elementary_information_and_technology 3.78 21.01 39.92 41.18 35.71 5.88 29.41
elementary_mathematics 4.35 12.17 31.74 33.91 24.35 4.35 17.39
ethnology 2.22 22.96 32.59 36.3 30.37 4.44 23.7
food_science 2.1 19.58 32.87 23.78 32.17 2.8 28.67
genetics 1.7 14.2 28.41 27.84 23.3 6.82 23.86
global_facts 0.67 18.12 33.56 30.87 38.26 7.38 19.46
high_school_biology 4.14 15.38 11.24 24.26 18.34 7.1 18.34
high_school_chemistry 0.76 9.09 13.64 10.61 14.39 10.61 19.7
high_school_geography 0.85 23.73 37.29 27.12 27.97 5.93 23.73
high_school_mathematics 5.49 11.59 31.71 25.0 33.54 5.49 20.12
high_school_physics 0.91 14.55 37.27 30.91 27.27 2.73 29.09
high_school_politics 2.1 18.18 26.57 18.88 28.67 5.59 21.68
human_sexuality 2.38 17.46 34.13 34.92 42.06 3.97 30.95
international_law 2.16 19.46 35.14 34.05 24.32 5.41 21.08
journalism 1.16 18.02 39.53 34.88 31.4 4.65 26.74
jurisprudence 1.95 17.52 35.77 27.98 36.25 8.03 21.65
legal_and_moral_basis 1.4 21.96 58.88 48.6 50.47 7.48 35.51
logical 2.44 15.45 36.59 27.64 39.02 4.07 26.02
machine_learning 0.82 10.66 23.77 27.05 20.49 4.1 21.31
management 2.86 18.1 40.0 38.1 32.86 5.71 24.76
marketing 2.22 20.0 43.89 42.78 37.22 6.11 31.11
marxist_theory 0.53 19.58 48.15 37.57 39.68 6.88 31.75
modern_chinese 3.45 18.97 20.69 21.55 18.97 3.45 12.93
nutrition 1.38 11.03 36.55 37.24 29.66 2.76 20.0
philosophy 2.86 18.1 44.76 37.14 40.0 5.71 28.57
professional_accounting 4.57 17.14 39.43 37.71 30.86 6.29 20.57
professional_law 0.95 16.11 26.07 18.96 25.12 4.74 17.06
professional_medicine 2.66 12.23 22.87 18.62 23.14 8.24 17.82
professional_psychology 2.59 18.97 36.64 34.05 35.78 4.74 25.0
public_relations 1.15 9.77 32.18 32.18 33.91 5.17 23.56
security_study 3.7 22.22 40.0 37.04 34.07 5.19 29.63
sociology 0.44 17.26 42.92 36.28 35.4 5.75 26.11
sports_science 2.42 22.42 33.33 33.33 30.3 7.27 24.85
traditional_chinese_medicine 1.62 15.68 27.03 27.57 24.32 4.86 17.3
virology 1.78 11.83 34.32 25.44 21.89 1.78 27.81
world_history 3.73 13.04 34.78 26.09 27.33 4.97 24.84
world_religions 2.5 11.88 45.62 31.87 35.0 2.5 21.25
Average 2.15 15.8 33.45 29.97 29.05 5.44 23.62
Table 16: Accuracy on 67 subjects of Ti-MMLU(DA)
Task Name Llama3.1-405B-Instruct Qwen2.5-7B-Instruct Qwen2.5-32B-Instruct Qwen2.5-72B-Instruct O1-mini Deepseek-r1
agronomy 25.44 0.0 13.02 7.69 5.33 20.71
anatomy 20.27 1.35 8.78 3.38 4.05 8.78
ancient_chinese 15.24 1.83 15.85 5.49 5.49 11.59
arts 23.75 3.12 19.38 8.75 4.38 17.5
astronomy 26.06 0.61 13.94 5.45 9.7 16.97
business_ethics 24.4 1.44 12.44 7.66 4.78 19.14
chinese_civil_service_exam 20.62 3.75 8.75 7.5 2.5 15.62
chinese_driving_rule 24.43 0.76 11.45 6.11 7.63 19.85
chinese_food_culture 24.26 0.0 13.97 7.35 2.94 18.38
chinese_foreign_policy 28.97 2.8 12.15 7.48 6.54 18.69
chinese_history 25.08 1.55 13.62 6.19 3.72 12.38
chinese_literature 21.08 1.96 16.18 9.31 6.86 19.61
chinese_teacher_qualification 34.08 1.12 15.64 8.94 5.03 24.58
clinical_knowledge 24.89 1.27 10.55 8.44 5.91 5.49
college_actuarial_science 16.04 3.77 13.21 8.49 6.6 24.53
college_education 39.25 2.8 16.82 9.35 10.28 20.56
college_engineering_hydrology 28.3 0.94 16.98 6.6 7.55 17.92
college_law 21.3 1.85 12.04 6.48 8.33 17.59
college_mathematics 18.1 6.67 7.62 9.52 2.86 7.62
college_medical_statistics 26.42 0.94 22.64 3.77 4.72 2.83
college_medicine 19.05 1.47 6.96 3.66 3.66 6.23
computer_science 28.43 0.49 11.27 4.41 4.41 11.27
computer_security 30.41 3.51 14.04 8.77 8.77 12.28
conceptual_physics 27.89 1.36 11.56 3.4 7.48 12.24
construction_project_management 23.74 1.44 10.07 5.76 5.04 17.27
economics 18.87 3.14 13.21 3.77 2.52 18.24
education 28.83 0.61 11.66 8.59 6.13 17.79
electrical_engineering 25.58 0.58 8.14 8.14 5.23 22.09
elementary_chinese 17.06 1.19 13.89 8.33 3.97 18.65
elementary_commonsense 28.28 2.02 15.66 5.56 3.54 24.24
elementary_information_and_technology 33.19 2.1 14.29 4.2 7.98 13.87
elementary_mathematics 28.26 6.09 20.0 7.83 4.78 37.39
ethnology 25.93 2.22 18.52 12.59 5.19 33.33
food_science 32.87 0.0 16.08 6.99 1.4 13.29
genetics 21.59 0.57 12.5 6.25 7.39 6.25
global_facts 20.81 1.34 16.11 8.05 6.71 14.09
high_school_biology 20.71 1.18 7.69 4.14 6.51 1.78
high_school_chemistry 19.7 1.52 3.79 4.55 6.82 3.03
high_school_geography 22.88 1.69 11.02 7.63 5.08 20.34
high_school_mathematics 24.39 17.68 22.56 13.41 6.71 23.78
high_school_physics 23.64 5.45 14.55 6.36 4.55 20.91
high_school_politics 26.57 2.8 14.69 3.5 2.1 20.98
human_sexuality 27.78 0.0 16.67 10.32 6.35 16.67
international_law 22.16 1.62 21.08 9.73 5.95 2.7
journalism 26.74 1.74 13.37 8.72 6.4 20.35
jurisprudence 27.01 1.95 12.41 9.25 7.06 17.52
legal_and_moral_basis 35.51 0.93 17.76 10.28 6.54 34.11
logical 30.08 2.44 17.07 5.69 8.13 12.2
machine_learning 24.59 0.82 8.2 3.28 3.28 10.66
management 24.29 0.48 13.81 7.14 11.43 21.43
marketing 31.11 0.56 18.89 10.0 11.67 16.67
marxist_theory 23.81 2.65 10.05 8.47 7.41 22.22
modern_chinese 19.83 1.72 12.07 2.59 6.03 14.66
nutrition 27.59 0.69 8.28 3.45 7.59 10.34
philosophy 27.62 1.9 16.19 7.62 5.71 0.95
professional_accounting 22.29 0.57 13.71 5.14 6.29 15.43
professional_law 18.48 0.0 9.95 5.69 4.74 7.58
professional_medicine 18.09 0.53 12.5 6.91 7.18 6.91
professional_psychology 21.12 0.86 18.53 6.47 6.9 16.81
public_relations 24.14 1.72 17.24 5.17 8.62 18.39
security_study 26.67 3.7 14.07 14.81 5.93 12.59
sociology 25.22 0.88 15.93 4.42 6.19 16.81
sports_science 34.55 0.61 21.21 10.91 7.88 13.33
traditional_chinese_medicine 21.08 1.08 10.81 7.03 7.03 9.73
virology 26.04 0.0 11.24 4.14 11.83 7.69
world_history 24.22 3.11 17.39 10.56 8.07 16.77
world_religions 27.5 0.62 13.12 7.5 3.12 27.5
Average 24.99 1.91 13.8 7.09 6.1 15.82
Table 17: Accuracy on 67 subjects of Ti-MMLU(DA) continue
Category Specific Content
China specific ancient_chinese
traditional_chinese_medicine
modern_chinese
high_school_politics
ethnology
elementary_commonsense
elementary_chinese
education
construction_project_management
chinese_teacher_qualification
chinese_literature
chinese_history
chinese_foreign_policy
chinese_food_culture
chinese_driving_rule
chinese_civil_service_exam
Other sports_science
professional_medicine
nutrition
legal_and_moral_basis
human_sexuality
food_science
elementary_it
computer_security
college_medicine
clinical_knowledge
agronomy
Social Sciences sociology
security_study
public_relations
professional_psychology
professional_accounting
marketing
management
journalism
high_school_geography
economics
college_education
business_ethics
Humanities world_religions
world_history
professional_law
philosophy
marxist_theory
logical
jurisprudence
international_law
global_facts
college_law
arts
STEM virology
machine_learning
high_school_physics
high_school_mathematics
high_school_chemistry
high_school_biology
genetics
elementary_mathematics
electrical_engineering
conceptual_physics
computer_science
college_medical_statistics
college_mathematics
college_engineering_hydrology
college_actuarial_science
astronomy
anatomy
Table 18: Ti-MMLU Category Mapping Table