### v3

# TLUE: A Tibetan Language Understanding Evaluation Benchmark

Abstract
Large language models have made tremendous progress in recent years, but low-resource languages, like Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of LLMs. To address this gap, we present a Tibetan Language Understanding Evaluation Benchmark, TLUE, which is also the first large-scale benchmark for measuring the proficiency of large language models in the Tibetan language. TLUE comprises two major components: a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and a safety benchmark encompassing 7 subdomains. Finally, we evaluate a diverse set of state-of-the-art LLMs. Experimental results demonstrate that most large language models perform below the random baseline, especially highlighting the considerable challenges they face in Tibetan language processing. TLUE provides a crucial foundation for advancing future research in Tibetan language understanding and highlights the importance of promoting greater inclusivity in the development of large language models.
TLUE: A Tibetan Language Understanding Evaluation Benchmark
[…]
1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019). However, many mainstream LLMs, like GPT OpenAI (2024); Achiam et al. (2023), Claude Anthropic (2024), Gemini DeepMind (2024), LlaMA Dubey et al. (2024), Qwen Team (2024) and DeepSeek DeepSeek-AI et al. (2025, 2024) largely overlook low-resource languages, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025). Tibetan is one of the languages included, spoken by over seven million people. It has unique grammatical features and limited annotated data of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023); Liu et al. (2025).
Current language understanding benchmarks, including GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), and BIG-bench Srivastava et al. (2022), focus on high-resource languages, leaving low-resource languages like Tibetan without tailored evaluation frameworks, including benchmarks from China Li et al. (2023); Zhang et al. (2023). Even existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). In this paper, we introduce the TLUE, a Tibetan Language Understanding Evaluation benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation. Furthermore, we selected several state-of-the-art large language models for training and evaluation, and the experimental results demonstrate that TLUE effectively exposes the limitations of current models in handling Tibetan language tasks. Most LLMs show significantly lower performance on TLUE compared to benchmarks in high-resource languages, particularly struggling with tasks requiring deep linguistic understanding and domain-specific knowledge. This highlights the urgent need for more inclusive pretraining data and architecture adaptations tailored to low-resource languages like Tibetan. Our findings validate TLUE as a robust tool for evaluating and guiding the development of more equitable language models.
The main contributions of our work are summarized as follows:
  • We developed and publicly released TLUE, the first large-scale Tibetan Language Understanding Evaluation benchmark. It is designed to fill the gap in existing resources by providing a comprehensive suite of tasks that span 67 knowledge-based subdomains and 7 safety-critical categories, addressing both general understanding and ethical alignment in Tibetan.
  • We evaluated several state-of-the-art large language models, including GPT-4O Achiam et al. (2023) and GPT-O1-mini OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5 DeepMind (2024), LlaMA-3.1 Dubey et al. (2024), Qwen-2.5 Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024) and DeepSeek-R1 DeepSeek-AI et al. (2025), to systematically assess their capabilities in both general Tibetan language understanding and safety-oriented tasks. These LLMs represent a diverse set of architectures and training strategies, enabling a broad comparison of their performance under low-resource conditions.
  • Our qualitative and quantitative analyses reveal substantial limitations in current large language models when applied to Tibetan. Most models performed significantly below expected baselines, even falling below random choice levels on certain tasks, particularly in domains requiring complex reasoning or cultural sensitivity. These findings underscore the urgent need for research into more inclusive, linguistically-aware model development tailored to low-resource languages.
2 Related Work
2.1 Language Understanding Benchmark
LLMs have been extensively evaluated on multilingual and domain-specific benchmarks, such as XTREME Hu et al. (2020), XGLUE Liang et al. (2020), and MASSIVE FitzGerald et al. (2022). However, these benchmarks predominantly focus on high-resource languages, with little attention given to low-resource languages like Tibetan. CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023) introduced large-scale Chinese language evaluations, but no equivalent benchmark existed for Tibetan.
2.2 Low Resource Language Evaluation
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane $\forall$ et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation.
2.3 Safety and Ethical Limitations in LLM
Recent work has also examined the safety and ethical risks of LLMs, particularly in multilingual and low-resource contexts. Studies such as SafetyBench Zhang et al. (2023) and HolisticEval Liang et al. (2023) assess biases, fairness, and robustness in AI systems. However, these benchmarks largely exclude Tibetan and other low-resource languages, leaving significant gaps in understanding how safety concerns manifest in such linguistic environments.
2.4 Motivation
Based on the current progress of LLM in minority languages, we propose the TLUE, filling the gap by providing a comprehensive Tibetan evaluation dataset across multiple domains and safety-related tasks. We referred to the two benchmarks, CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), and asked ancient Tibetan experts to manually translate and verify them. In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE. TLUE can not only enable systematic evaluation of model performance in a challenging low-resource setting, but also introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
By offering a comprehensive, multi-domain framework for both language understanding and safety evaluation, TLUE fills a critical gap in the current landscape of LLM benchmarks, enabling systematic assessment of Tibetan language models in a challenging low-resource setting.
Figure 1: The Curation Process of TLUE
3 TLUE
In designing TLUE, we adhere to several key principles to ensure a rigorous and meaningful evaluation of Tibetan LLMs.
First, we align our benchmark with existing multilingual evaluation frameworks to facilitate direct comparison with established benchmarks. This allows researchers to assess the performance of Tibetan LLMs relative to models trained in other languages, providing a broader perspective on cross-lingual capabilities Park et al. (2024).
Additionally, we ensure that the evaluation data is carefully curated and free from data leakage, maintaining the integrity of benchmark results and preventing unintended memorization effects Doe and Smith (2024).
3.1 Curation Process
As shown in Figure 1, the TLUE benchmark follows a structured multi-stage process to ensure the quality and cultural relevance of Tibetan evaluation data in the curation process. Source datasets, such as CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), are first translated into Tibetan using Claude-3.5-Sonnet Anthropic (2024) , which demonstrates strong Tibetan text generation capabilities. The translations undergo rule-based filtering to correct syntactic inconsistencies and structural errors.
Following this, expert human annotators, including 2 Tibetan language specialists and their team of 5 annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
The final phase includes domain-specific validation, where legal, medical, and scientific content is reviewed by subject-matter experts for accuracy. This multi-layered curation approach ensures TLUE remains a comprehensive and culturally adapted benchmark for Tibetan language model evaluation.
3.2 Size
TLUE consists of 22,963 evaluation questions, divided into Ti-MMLU for knowledge assessment and Ti-SafetyBench for safety evaluation.
Ti-MMLU includes 11,528 multiple-choice questions across 67 subjects, covering social sciences, STEM, China-specific topics, humanities, and other domains, enabling a comprehensive evaluation of Tibetan language understanding. Ti-SafetyBench contains 11,435 questions across 7 safety categories, including ethics, bias, health, and privacy, designed to assess LLMs’ ability to handle sensitive content.
While Ti-MMLU focuses on broad, multi-domain knowledge evaluation, Ti-SafetyBench targets high-precision safety assessments, together forming a balanced and comprehensive Tibetan language benchmark. The dataset sizes are summarized in Appendix A Figure 5, which illustrates the comprehensive breakdown of Ti-MMLU and Ti-SafetyBench, highlighting the wide range of subjects and safety categories covered. For more details of such two sub-benchmarks, they are shown in Appendix A.
3.3 Why Extend Existing Benchmarks?
Developing a benchmark for Tibetan presents challenges such as limited annotated corpora and linguistic expertise. TLUE extends existing benchmarks via translation, enhanced with human annotation and quality control. Key reasons include:
3.3.1 Enabling Cross-Lingual Comparability
A translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
3.3.2 Addressing Data Scarcity and Leveraging Established Designs
Due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
3.3.3 Ensuring Quality through Controlled Adaptation
TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards. This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
4 Implementation
Our experiments cover a diverse range of LLMs, including both open-source and proprietary LLMs.
For Open-source LLM, we evaluate several open-source LLMs, including LlaMA-3.1-405B Dubey et al. (2024), LlaMA-3.1-8B Dubey et al. (2024), Qwen-2.5-72B Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These LLMs represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
For Proprietary LLM, we also evaluate several proprietary LLMs, including GPT-4O OpenAI (2024), GPT-3.5-Turbo Achiam et al. (2023), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
4.1 Experimental Setup
We investigate several key aspects of LLM performance in Tibetan:
$\bullet$ We evaluate all LLMs on Ti-MMLU and Ti-SafetyBench, measuring their Tibetan multi-task understanding and safety capabilities.
$\bullet$ We assess the impact of language resource availability by comparing CMMLU Li et al. (2023) and Ti-MMLU and extend this analysis to safety evaluation by comparing SafetyBench Zhang et al. (2023) with Ti-SafetyBench.
$\bullet$ We explore the differences between reasoning-optimized LLMs and chat LLMs, examining whether reasoning enhancements improve low-resource language performance.
$\bullet$ We study the effect of LLM size on Tibetan language understanding, comparing different parameter variants of LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
4.2 Evaluation
4.2.1 Evaluation Methodology
We employ zero-shot evaluation to simulate real-world conditions where models must perform tasks without task-specific examples, providing a more accurate assessment of performance in Tibetan. Appendix Figure 7 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
4.2.2 Evaluation Metrics
We evaluate LLM performance on TLUE using two metrics: response rate and accuracy. Response Rate measures the proportion of valid responses, while Accuracy tracks the proportion of correct answers. We also compute Conditional Accuracy, which reflects the LLM’s performance when giving valid responses.
For both Ti-MMLU and Ti-SafetyBench, We use two methods for metric calculation:
  • Direct Answer Calculation measures if the LLM selects a valid and correct answer.
  • Concern All Answer Calculation evaluates the LLM’s maximum performance by considering all valid options and selecting correct answers based on remaining choices.
These methods provide a thorough assessment of LLMs’ Tibetan language performance. For more details of Ti-MMLU and Ti-SafetyBench, please refer to Appendix B Algorithm 1 and Appendix B Algorithm 2.
5 Performance on TLUE
This section shows the performance of LLM on the TLUE benchmark. We use abbreviated forms for some of the special names or evaluation matrix names. Tables 1(a) and 1(b) show results on Ti-MMLU and Ti-SafetyBench, evaluating Tibetan language understanding and safety.
Method LLM Version Avg. STEM Human Social Other China
DA Claude 3.5-Sonnet 33.95 28.04 36.32 39.31 34.01 32.09
Gemini 1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7
DeepSeek R1 15.74 13.87 14.24 18.06 13.94 18.58
DeepSeek V3 29.51 23.57 31.97 33.65 29.92 28.44
GPT 4O 16.00 12.73 16.36 18.04 16.92 15.96
GPT 3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87
GPT O1-mini 6.14 6.15 6.33 7.17 6.06 4.98
Qwen 2.5-32B 13.94 12.63 14.98 14.92 13.71 13.44
Qwen 2.5-72B 7.27 6.07 7.98 7.52 7.74 7.02
Qwen 2.5-7B 1.8 2.94 1.87 1.63 0.9 1.68
LlaMA 3.1-405B 25.08 23.88 24.25 25.58 27.62 24.07
LlaMA 3.1-70B 23.73 23.16 23.2 26.2 24.65 21.45
LlaMA 3.1-8B 5.47 5.48 5.56 5.99 5.46 4.86
CAA Claude 3.5-Sonnet 35.63 30.88 37.47 40.58 35.26 33.96
Gemini 1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81
DeepSeek R1 27.45 21.01 25.99 32.72 26.44 31.08
DeepSeek V3 32.16 27.03 34.58 36.26 32.00 30.94
GPT 4O 17.51 14.25 17.71 19.69 18.46 17.45
GPT 3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16
GPT O1-mini 9.67 9.69 9.80 10.14 9.68 9.02
Qwen 2.5-32B 18.56 16.66 20.3 19.72 17.47 18.67
Qwen 2.5-72B 16.50 15.73 17.88 17.00 15.84 16.04
Qwen 2.5-7B 14.59 13.92 13.66 16.34 14.57 14.46
LlaMA 3.1-405B 25.28 24.10 24.50 25.87 27.73 24.22
LlaMA 3.1-70B 23.79 23.22 23.24 26.31 24.65 21.52
LlaMA 3.1-8B 7.44 7.95 7.54 7.38 7.41 6.92
- Random - 25.00 25.00 25.00 25.00 25.00 25.00
(a) Accuracy Performance on the Ti-MMLU
Method LLM Version Avg. OFF UB PH MH IA EM PP
DA Claude 3.5-Sonnet 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini 1.5-Flash 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
DeepSeek R1 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
DeepSeek V3 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
GPT 4O 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT 3.5-turbo 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
GPT O1-mini 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Qwen 2.5-72B 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen 2.5-32B 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen 2.5-7B 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
LlaMA 3.1-405B 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
LlaMA 3.1-70B 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
LlaMA 3.1-8B 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA Claude 3.5-Sonnet 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini 1.5-Flash 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
DeepSeek R1 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
DeepSeek V3 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
GPT 4O 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT 3.5-turbo 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
GPT O1-mini 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Qwen 2.5-72B 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen 2.5-32B 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen 2.5-7B 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
LlaMA 3.1-405B 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
LlaMA 3.1-70B 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
LlaMA 3.1-8B 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
- Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
(b) Accuracy Performance on the Ti-SafetyBench
Table 1: Accuracy Performance of LLMs on the TLUE based on CAA and DA ($\times$100%)
5.1 Performance on Ti-MMLU
As shown in Table 1(a), most LLMs perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary LLMs, the GPT Achiam et al. (2023); OpenAI (2024) series underperforms, whereas open-source models like DeepSeek-V3 DeepSeek-AI et al. (2024) exceed the baseline. For reasoning-optimized LLMs, DeepSeek-R1 DeepSeek-AI et al. (2025) surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
The performance gap between proprietary and open-source models is minimal, with DeepSeek-V3 DeepSeek-AI et al. (2024) and Gemini-1.5-Flash DeepMind (2024) performing comparably to Claude-3.5-Sonnet Anthropic (2024). Notably, GPT-3.5-Turbo Achiam et al. (2023); OpenAI (2024) underperforms relative to LlaMA-3.1-8B Dubey et al. (2024), suggesting that pretraining data, adaptation, and optimization significantly impact Tibetan language performance.
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, LlaMA-3.1-405B Dubey et al. (2024) excels in Other, and Qwen-2.5-72b Team (2024) in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
The detial data is released in Appendix Table 15 and Appendix Table 16.
5.2 Performance on Ti-SafetyBench
As shown in Table 1(b), most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy (58.5% in CAA), whereas GPT-4O Achiam et al. (2023) and GPT-3.5-Turbo Achiam et al. (2023) underperform, with the latter significantly below random.
Among open-source models, DeepSeek-V3 DeepSeek-AI et al. (2024) surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. LlaMA-3.1-405B Dubey et al. (2024) approaches the baseline, while smaller models like LlaMA-3.1-8B Dubey et al. (2024) and Qwen-2.5-72B Team (2024) perform poorly. Reasoning-optimized models such as DeepSeek-R1 DeepSeek-AI et al. (2025) excel in CAA but struggle in DA, whereas O1-mini OpenAI (2024) consistently underperforms.
These findings suggest that safety alignment in Tibetan remains challenging, with pretraining data, adaptation strategies, and model architecture playing a more critical role than accessibility.
5.3 High-Resource vs. Low-Resource
To examine the performance disparity between high-resource and low-resource languages, we compare model accuracy on CMMLU Li et al. (2023) and Ti-MMLU using the CAA evaluation. Figure 2 illustrates performance differences across domains, while Appendix Table 6 presents detailed accuracy comparisons.
5.3.1 Performance Degradation
All LLMs experience substantial accuracy drops from CMMLU Li et al. (2023) to Ti-MMLU. Qwen-2.5-72B Team (2024) declines from 84.70% $\rightarrow$ 16.50%, GPT-4 Achiam et al. (2023) from 68.90% $\rightarrow$ 17.51%, and ChatGPT Achiam et al. (2023); OpenAI (2024) from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%), underscoring the difficulty of adapting LLMs to low-resource languages.
5.3.2 Domain Performance Shifts
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU Li et al. (2023), LLMs excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. For instance, Qwen-2.5-72B Team (2024) achieves the highest accuracy in "Other" (87.35%) on CMMLU Li et al. (2023) but shifts to "Humanities" (17.88%) on Ti-MMLU. LlaMA-3.1-70B Dubey et al. (2024) transitions from "Other" (74.72%) in CMMLU Li et al. (2023) to "Social Sciences" (26.31%) in Ti-MMLU. GPT-4’s Achiam et al. (2023) best-performing category changes from "Other" (73.16%) in CMMLU Li et al. (2023) to "Social Sciences" (19.69%) in Ti-MMLU (GPT-4O Achiam et al. (2023)). These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
5.3.3 STEM as the Weakest Domain
STEM remains the most challenging domain in both benchmarks. It consistently ranks lowest in CMMLU Li et al. (2023) and Ti-MMLU, except for LlaMA-3.1-70B Dubey et al. (2024), which performs worst in "China-specific" tasks. This indicates significant difficulties in mathematical and technical reasoning in Tibetan.
5.3.4 Ranking Inconsistencies
LLM rankings in CMMLU Li et al. (2023) do not consistently translate to Ti-MMLU. For instance, LlaMA-3.1-70B Dubey et al. (2024) ranks below Qwen-2.5-72B Team (2024) in CMMLU Li et al. (2023) but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
5.3.5 Implication: Low-Resource Adaptation
The substantial performance gap underscores the need for improved Tibetan data coverage in pretraining and fine-tuning. While LLMs excel in high-resource languages, their struggles in Tibetan highlight the critical role of data availability and adaptation strategies in enhancing low-resource language understanding.
5.4 Safety in Low-Resource Languages
(a) Performance Comparison of Qwen-2.5-72B Team (2024) from CMMLU Li et al. (2023) to Ti-MMLU
(b) Performance Comparison of LlaMA-3.1-70B Dubey et al. (2024) from CMMLU Li et al. (2023) to Ti-MMLU
(c) Performance Comparison of GPT-4o Achiam et al. (2023) from CMMLU Li et al. (2023) to Ti-MMLU
(d) Performance Comparison of GPT-3.5-turbo Achiam et al. (2023) from CMMLU Li et al. (2023) to Ti-MMLU
Figure 2: LLMs’ Performance Degradation from CMMLU Li et al. (2023) to Ti-MMLU
(a) Performance Comparison of GPT-4o Achiam et al. (2023) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
(b) Performance Comparison of GPT-3.5-turbo Achiam et al. (2023) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
(c) Performance Comparison of Qwen-2.5 (Chat) Team (2024) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
Figure 3: LLMs’ Performance Degradation from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
To evaluate LLMs’ safety in low-resource languages, we compare performance on SafetyBench Zhang et al. (2023) and Ti-SafetyBench using CAA. Figure 3 illustrates accuracy drops across safety categories, while Appendix Table 7 provides detailed performance breakdowns. These results highlight the challenges of maintaining safety alignment in Tibetan.
5.4.1 Performance Degradation in Tibetan
LLMs show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 Achiam et al. (2023) achieves 89.2% on SafetyBench Zhang et al. (2023), but GPT-4O Achiam et al. (2023), despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo Achiam et al. (2023) experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B Team (2024) decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
5.4.2 Variability in Safety Categories
Performance degradation varies across safety categories. GPT-4 Achiam et al. (2023) excels in PH and EM on SafetyBench Zhang et al. (2023), whereas GPT-4O Achiam et al. (2023), despite being more advanced, performs best in IA on Ti-SafetyBench, highlighting the challenges of transferring safety alignment to Tibetan.
Similarly, Qwen-chat-7B Team (2024) outperforms random in EM and IA on SafetyBench Zhang et al. (2023), but Qwen-2.5-7B Team (2024) retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
GPT-3.5-Turbo Achiam et al. (2023) falls well below the random baseline on Ti-SafetyBench, while GPT-4O Achiam et al. (2023) and Qwen-2.5-7B Team (2024) exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
5.4.3 Low-Resource Safety Alignment
The sharp decline in performance on Ti-SafetyBench underscores the challenge of maintaining safety alignment in Tibetan. While models perform well on SafetyBench Zhang et al. (2023), most fall below or barely match the baseline on Ti-SafetyBench. Even top-performing models struggle with safety generalization, highlighting the need for targeted fine-tuning and enhanced multilingual adaptation in low-resource environments.
5.5 Reasoning vs. Chat
We compare reasoning-optimized and chat LLMs on TLUE, analyzing response behavior, accuracy, and conditional accuracy. The results are summarized in Appendix C Table 8, Appendix C Table 9 and Appendix C Table 12, with a broader comparison provided on Ti-MMLU (Appendix C Table 10) and Ti-SafetyBench (Appendix C Table 11), covering overall accuracy across knowledge and safety domains.
5.5.1 Response Rate and Output Behavior
Reasoning LLMs, compared to chat LLMs, tend to evaluate all answer choices before selecting the correct one, resulting in higher response rates for CAA than DA (Appendix C Table 8 and Appendix C Table 9). This behavior stems from a step-by-step evaluation process. When evaluating reasoning models, intermediate reasoning steps are filtered, and only the final output is considered, as outlined in Appendix B Algorithm 3 and Appendix B Algorithm 4.
5.5.2 Impact: Response Rate on Performance
Lower response rates notably affect reasoning models’ performance, highlighting the inherent difficulty in responding to Tibetan prompts. The gap between DA and CAA accuracy underscores this challenge, with DA tasks requiring direct answer generation (Appendix C Table 8 and Appendix C Table 9). Reasoning models, unlike chat models, struggle more with Tibetan prompts due to their emphasis on logical inference over language generation.
5.5.3 Generalization of Reasoning Models in Low-Resource Languages
As shown in Appendix C Table 12, DeepSeek-R1 Ranathunga et al. (2023) consistently outperforms DeepSeek-V3 DeepSeek-AI et al. (2024) in conditional accuracy. Additionally, O1-mini OpenAI (2024) outperforms GPT-4O Achiam et al. (2023) in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini OpenAI (2024) underperforms compared to GPT-4O Achiam et al. (2023) in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
5.6 Model Scale and Low-Resource Performance
We investigate the effect of model scale on Tibetan language understanding and safety alignment using Ti-MMLU (Figure 4(a) , Appendix C Table 13) and Ti-SafetyBench (Figure 4(b), Appendix C Table 14).
5.6.1 Effect: Scale on Tibetan Understanding
LLMs generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LlaMA-3.1-405B Dubey et al. (2024) slightly outperforms LlaMA-3.1-70B Dubey et al. (2024), while Qwen-2.5-32B Team (2024) surpasses Qwen-2.5-72B Team (2024). Notably, Qwen-2.5-7B Team (2024) achieves comparable or superior accuracy to Qwen-2.5-72B Team (2024) in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
5.6.2 Impact: Scale on Safety Alignment
LLMs show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen-2.5-72B Team (2024) significantly outperforms Qwen-2.5-32B Team (2024), indicating that scaling benefits safety alignment more than general understanding. In contrast, LlaMA-3.1-70B Dubey et al. (2024) shows only marginal improvement over LlaMA-3.1-8B Dubey et al. (2024), suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
5.6.3 Challenge: Scaling for Low-Resource
While LLMs generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller LLMs, such as Qwen-2.5-7B Team (2024), achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.
(a) Average accuracy on Ti-MMLU
(b) Average accuracy on Ti-SafetyBench
Figure 4: Average accuracy on TLUE across different model scales for LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
6 Conclusion
This paper presents TLUE, a benchmark for evaluating large language models in Tibetan, covering general knowledge understanding (Ti-MMLU) and safety alignment (Ti-SafetyBench). We show that both proprietary and open-source models struggle in Tibetan, performing below or near the random baseline. We identify key challenges, including significant performance drops when transitioning from high-resource languages to Tibetan, limitations in current multilingual training, and the need for enhanced Tibetan language resources. While open-source models show promising domain-specific improvements, reasoning-optimized models demonstrate better generalization but face difficulties in following Tibetan prompts. Model scaling provides inconsistent benefits, with architecture, data quality, and fine-tuning strategies proving more crucial than model size. These results underscore the need for improved pretraining and targeted fine-tuning, with TLUE serving as a foundation for future low-resource language modeling research, promoting inclusivity and robustness in LLM development.
7 Limitation
The limitation of TLUE is the limited coverage of Tibetan cultural and folk knowledge. While the benchmark includes diverse linguistic and domain-specific tasks, it lacks evaluation datasets tailored to Tibetan folklore, traditional customs, and region-specific knowledge. As cultural understanding is crucial for real-world applications, future work should expand TLUE to better assess LLMs’ capabilities in these areas.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.
Appendix A Appendix: Details of TLUE
As shown in Figure 5, TLUE consists of 2 sub-benchmarks: Ti-MMLU and Ti-SafetyBench. Together, these two components offer a comprehensive evaluation of both the general knowledge proficiency and safety alignment of LLMs in Tibetan. TLUE emphasizes challenges unique to low-resource languages and supports zero-shot and few-shot settings to facilitate cross-model comparisons without relying on language-specific tuning.
Figure 6: Statistical Categories of the Ti-MMLU Benchmark
A.1 Ti-MMLU
Ti-MMLU [^1] License: CC-BY-NC-SA 4.0 is an evaluation sub-benchmark designed specifically for Tibetan LLMs, similar to MMLU Hendrycks et al. (2020) in the English field and CMMLU Li et al. (2023) in the Chinese field. It comprehensively tests the LLM’s knowledge understanding and reasoning capabilities in a multi-disciplinary and multi-task environment through multiple-choice questions.
As shown in Firgure 6, Ti-MMLU contains 67 subtasks, covering multiple subject areas from middle school to university and even professional examinations, such as mathematics, physics, history, law, medicine, engineering, philosophy, literature, etc., covering the unique local knowledge system in Tibetan areas, such as college entrance examinations, teacher qualification certificates, medical examinations, etc., and is particularly suitable for evaluating the LLM’s mastery of Tibetan language context and professional knowledge.
Ti-MMLU uses zero-shot or few-shot settings, does not provide contextual learning, and directly examines the generality and true capabilities of the model. It is not only suitable for model comparison and ranking, but also helps developers discover the weak links of the model in specific fields, such as law and medicine.
Figure 7: Prompt and Example of Ti-MMLU
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 8: Prompt and Example of Ti-SafetyBench
One sample is shown in Figure 7.
A.2 Ti-SafetyBench
Ti-SafetyBench [^2] License: Apache License 2.0 is a multilingual benchmark systematically evaluate the safety of LLMs, similar to SafetyBench Zhang et al. (2023), when handling sensitive or high-risk Tibetan content. It consists of multiple-choice questions, and supports zero-shot and few-shot evaluation settings to enable standardized comparisons across different models.
Ti-SafetyBench covers eight core safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, Privacy and Property, and Safety-Related Reasoning. These categories reflect common areas of concern in real-world applications where unsafe or harmful responses from LLMs may occur.
By providing a structured, quantifiable framework for assessment, Ti-SafetyBench serves as a critical tool for developers, researchers, and policymakers in improving and regulating the deployment of safe and trustworthy AI systems for Tibetan.
One sample is shown in Figure 8.
A.3 Human Evaluation
2 Tibetan language specialists and their team of 5 annotators refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. They are authors of this paper.
In the early days, we used Google Translate and Claude-3.5-Sonnet Anthropic (2024) for translation, which was then optimized by our experts. As shown in Table 2, the translation quality is poor, whether it is manually evaluated or using the BLEU matrix Papineni et al. (2002).
Evaluation Matrix Google Translate Claude-3.5-Sonnet
Expert Approval Rate 11.54% 28.74%
Domain Knowledge Alignment Score 0.95 2.3
Cultural Alignment Score 0.85 1.9
BLEU 23.2 34.8
Table 2: Comparison of Translation Quality Between Google Translate and Claude-3.5-Sonnet
Final translations were further refined and verified by Tibetan experts through multiple rounds of review, as described in our paper.
As shown in Table 3, the original Claude 3.5-translated dataset had an expert approval rate of 28.74%, which increased to 82.33% after the first expert alignment pass, and reached 100% following the second round of refinement. Corresponding domain and cultural alignment scores improved from 2.3/1.9 to 4.6/4.4, respectively. These results demonstrate the effectiveness of our alignment pipeline in reducing translationese artifacts and producing high-quality, culturally appropriate Tibetan-language content.
Stage Expert Approval Rate Domain Alignment Score Cultural Alignment Score
Initial 28.74% 2.3 1.9
1st Alignment 82.33% 3.7 3.5
2nd Alignment 100% 4.6 4.4
Table 3: Comparison of Translation Quality Between Google Translate and Claude-3.5-Sonnet
Specifically, we implemented a two-stage human refinement process following LLM-based translation, focused on both domain knowledge alignment and cultural alignment. This process was led by two Tibetan language experts and supported by a team of 5 trained annotators. Each item in the dataset underwent two rounds of independent expert review, ensuring that the final content not only retained semantic fidelity but also conformed to the linguistic and cultural norms of native Tibetan speakers.
We will also include an example (Figure 9) of expert annotation to illustrate how domain and cultural considerations were incorporated during the alignment process.
Figure 9: One Example of Expert Alignment
Appendix B Appendix: Algorithm of DA & CCA
Algorithm 1 describes the Direct Answer extraction process, which aims to identify a single valid choice (e.g., A/B/C/D) from a model’s output. It first filters out irrelevant content based on a user-defined exclusion list. Then, it scans for valid choices defined by the option count (2/3/4). If exactly one valid choice is found, it is returned as the model’s final answer; otherwise, the answer is considered invalid.
Algorithm 2 defines the Concern All Answer extraction strategy, which identifies a valid answer while tolerating noisy or repeated outputs. It first collects all valid choices in order, then removes full unordered sets of all options (e.g., ABCD) to eliminate exhaustive guessing behavior. If a single unique choice remains after filtering, it is returned; otherwise, no answer is selected.
Algorithm 3 presents a reasoning-aware version of Direct Answer extraction. To account for multi-step outputs from reasoning-oriented LLMs, model-specific reasoning content (e.g., Reasoning…Reasoned, <think>…</think>) is first removed. The algorithm then searches for a single valid answer choice (A/B/C/D) in the cleaned output. If exactly one valid choice is found, it is returned as the final prediction.
Algorithm 4 extends the Concern All Answer extraction method by incorporating reasoning content filtering for step-by-step models. It first removes model-specific reasoning traces, then collects all valid answer choices while discarding full unordered choice sets (e.g., ABCD). If a single unique choice remains after filtering, it is returned as the prediction; otherwise, the answer is deemed invalid.
Algorithm 1 Algorithm for Direct Answer Calculation (DA) with Variable Choice Count
1:procedure Extract Direct Answer(model_result, exclude_list, option_num)
2:     $valid\_choices\_map\leftarrow\{\text{`2''}:\{A,B\},\text{3''}:\{A,B,C\},\text{4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
9:     end for
10:     $found\_choices\leftarrow[choice\in valid\_choices\mid choice\in model\_result]$
11:     if $|found\_choices|=1$ then
12:         return $found\_choices[0]$triangleright Return extracted answer if unique
13:     else
14:         return $\emptyset$ $\triangleright$ Return empty if no valid answer is found
15:     end if
16:end procedure
Algorithm 2 Algorithm for Concern All Answer Calculation (CAA) with Variable Choice Count
1:procedure ExtractCAAAnswer(model_result, exclude_list, option_num)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options
7:     for each $excluded$ in $exclude\_list$ do
8:         Remove $excluded$ from $model\_result$
9:     end for
10:     Step 1: Extract all valid choices in order
11:     $all\_choices\leftarrow[ch\in model\_result\mid ch\in valid\_choices]$
12:     Step 2: Remove full valid choice combinations (unordered)
13:     $filtered\_choices\leftarrow[]$, $buffer\leftarrow[]$
14:     for each $ch$ in $all\_choices$ do
15:         Append $ch$ to $buffer$
16:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
17:              Clear $buffer$triangleright Remove one full valid choice combination
18:         else if $|buffer|>option\_num$ then
19:              Append first elements of $buffer$ to $filtered\_choices$
20:              Keep only last $(option\_num-1)$ elements in $buffer$
21:         end if
22:     end for
23:     Append remaining elements in $buffer$ to $filtered\_choices$
24:     Step 3: Determine final answer
25:     $unique\_choices\leftarrow set(filtered\_choices)$
26:     if $|unique\_choices|=1$ then
27:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains
28:     else
29:         return $\emptyset$ $\triangleright$ Return empty if multiple choices remain
30:     end if
31:end procedure
Algorithm 3 Algorithm for Direct Answer Extraction with Reasoning Filter
1:procedure ExtractDirectAnswer(model_result, exclude_list, option_num, model_name)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Filter model-specific reasoning content
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “DeepSeek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
11:     end if
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
14:     end for
15:     $found\_choices\leftarrow[choice\in valid\_choices\mid choice\in model\_result]$
16:     if $|found\_choices|=1$ then
17:         return $found\_choices[0]$triangleright Return extracted answer if unique
18:     else
19:         return $\emptyset$ $\triangleright$ Return empty if no valid answer is found
20:     end if
21:end procedure
Algorithm 4 Algorithm for Concern All Answer Calculation (CAA) with Reasoning Filter
1:procedure ExtractCAAAnswer(model_result, exclude_list, option_num, model_name)
2:     $valid\_choices\_map\leftarrow\{\text{``2''}:\{A,B\},\text{``3''}:\{A,B,C\},\text{``4''}:\{A,B,C,D\}\}$
3:     if $option\_num\notin valid\_choices\_map$ then
4:         return $\emptyset$ $\triangleright$ Invalid option number
5:     end if
6:     $valid\_choices\leftarrow valid\_choices\_map[option\_num]$triangleright Set of valid answer options $\triangleright$ Apply model-specific reasoning content filtering
7:     if $model\_name$ contains “o1” then
8:         Remove text matching ‘‘Reasoning.*?Reasoned .*? seconds’’ from $model\_result$
9:     else if $model\_name$ contains “DeepSeek-r1” then
10:         Remove text enclosed within ‘‘<think>...</think>’’ from $model\_result$
11:     end if
12:     for each $excluded$ in $exclude\_list$ do
13:         Remove $excluded$ from $model\_result$
14:     end for
15:     Step 1: Extract all valid choices in order
16:     $all\_choices\leftarrow[ch\in model\_result\mid ch\in valid\_choices]$
17:     Step 2: Remove full valid choice combinations (unordered)
18:     $filtered\_choices\leftarrow[]$, $buffer\leftarrow[]$
19:     for each $ch$ in $all\_choices$ do
20:         Append $ch$ to $buffer$
21:         if $|buffer|=option\_num$ and $set(buffer)=valid\_choices$ then
22:              Clear $buffer$triangleright Remove one full valid choice combination
23:         else if $|buffer|>option\_num$ then
24:              Append first elements of $buffer$ to $filtered\_choices$
25:              Keep only last $(option\_num-1)$ elements in $buffer$
26:         end if
27:     end for
28:     Append remaining elements in $buffer$ to $filtered\_choices$
29:     Step 3: Determine final answer
30:     $unique\_choices\leftarrow set(filtered\_choices)$
31:     if $|unique\_choices|=1$ then
32:         return $list(unique\_choices)$triangleright Return if exactly one unique choice remains
33:     else
34:         return $\emptyset$ $\triangleright$ Return empty if multiple choices remain
35:     end if
36:end procedure
Appendix C Appendix: Extended Experiment
For clarity and conciseness of TLUE, some indicators of Ti-MMLU and Ti-SafetyBench are presented as abbreviations in Table 4.
Full Name Abbreviation
Average Avg.
STEM STEM
Humanities Human
Social Sciences Social
Other Other
China Specific China
Offensiveness OFF
Unfairness and Bias UB
Physical Health PH
Mental Health MH
Illegal Activities IA
Ethics and Morality EM
Privacy and Property PP
Response Rate RR
Accuracy ACC
Conditional Accuracy CA
Direct Answer Calculation DA
Concern All Answer Calculation CAA
Table 4: List of Abbreviations for Professional Terms
C.1 Parameters Settings of LLMs
As shown in Table 5, regarding open-source model execution, the Qwen-2.5 Team (2024) and DeepSeek families DeepSeek-AI et al. (2024, 2025) were evaluated via their respective official APIs provided by the model developers. The LlaMA-3.1 Dubey et al. (2024) (8B, 70B, 405B) were accessed and run using the LlaMA-API platform. These details will also be clearly documented in the final version to enhance transparency and reproducibility.
LLM Version Temperature Top_p Stream
Claude 3-5-sonnet 1.0 None False
Gemini 1.5-flash None 0.95 False
GPT 3.5-turbo 1.0 1.0 False
4O 1.0 1.0 False
O1-mini 1.0 1.0 True
DeepSeek V3 1.0 None False
R1 1.0 None Rtue
LlaMA 3.1-8B 0.6 0.9 False
3.1-70B 0.6 0.9 False
3.1-405B 0.6 0.9 False
Qwen 2.5-7b 0.7 0.8 False
2.5-32b 0.7 0.8 False
2.5-72b 0.7 0.8 False
Table 5: Hyperparameters of LLM
C.2 Experimental Details
Table 6 and Table 7 shows a significant performance drop in Ti-MMLU and Ti-SafetyBench, with most results falling below the random baseline, highlighting the difficulty of safety alignment in low-resource languages.
LLM Version Benchmark Avg. STEM Human Social Other China
Qwen 2.5-72B CMMLU 84.70 80.67 87.00 84.66 87.35 83.21
Ti-CMMLU 16.50 15.73 17.88 17.00 15.84 16.04
LlaMA 3.1-70B CMMLU 69.01 61.60 71.44 69.42 74.72 63.79
Ti-CMMLU 23.79 23.22 23.24 26.31 24.65 21.52
GPT 4O CMMLU 68.90 63.16 69.19 70.26 73.16 63.47
Ti-CMMLU 17.51 14.25 17.71 19.69 18.46 17.45
3.5-Turbo CMMLU 53.22 44.80 53.61 54.22 59.95 49.74
Ti-CMMLU 3.42 3.82 3.35 3.68 3.09 3.16
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 6: Comparison of LLMs Performance on CMMLU Li et al. (2023) and Ti-MMLU ($\times$100%)
LLM Version Benchmark Avg. OFF UB PH MH IA EM PP
GPT 4O SafetyBench 89.2 85.4 76.4 95.5 94.1 92.5 92.6 92.5
Ti-SafetyBench 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
3.5-Turbo SafetyBench 80.4 76.1 68.7 78.4 89.7 87.3 78.5 87.9
Ti-SafetyBench 11.6 12.5 16 11.4 10.3 8.2 10.8 11.8
Qwen 2.5-7B SafetyBench 77.4 72.4 64.4 71.5 89.3 84.9 78.2 82.4
Ti-SafetyBench 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 7: Performance Comparison of LLMs on SafetyBench Zhang et al. (2023) and Ti-SafetyBench ($\times$100%)
LLM Version Method RR ACC CA
DeepSeek V3 DA 76.99 29.51 38.30
CAA 84.73 32.16 37.90
R1 DA 34.84 15.74 43.75
CAA 57.45 27.45 47.01
GPT 4O DA 51.68 16.00 30.93
CAA 57.47 17.51 30.44
O1-mini DA 22.71 6.14 27.4
CAA 35.00 9.67 27.74
Table 8: Response Rate, Accuracy, and Conditional Accuracy for Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method RR ACC
DeepSeek V3 DA 61.34 37.4
CAA 78.51 48.3
R1 DA 40.80 24.3
CAA 71.86 46.8
GPT 4O DA 61.41 31.1
CAA 65.45 32.9
O1-mini DA 28.48 10.9
CAA 38.31 15.1
Table 9: Response Rate and Accuracy for Reasoning-optimized and Chat LLMs on Ti-SafetyBench ($\times$100%)
LLM Version Method Avg. STEM Human Social Other China
DeepSeek V3 DA 29.51 23.57 31.97 33.65 29.92 28.44
CAA 32.16 27.03 34.58 36.26 32.00 30.94
R1 DA 15.74 13.87 14.24 18.06 13.94 18.58
CAA 27.45 21.01 25.99 32.72 26.44 31.08
GPT 4O DA 16.00 12.73 16.36 18.04 16.92 15.96
CAA 17.51 14.25 17.71 19.69 18.46 17.45
O1-mini DA 6.14 6.15 6.33 7.17 6.06 4.98
CAA 9.67 9.69 9.80 10.14 9.68 9.02
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 10: Accuracy Comparison of Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method Avg. OFF UB PH MH IA EM PP
DeepSeek V3 DA 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
CAA 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
R1 DA 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
CAA 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
GPT 4O DA 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
CAA 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
O1-mini DA 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
CAA 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 11: Accuracy Comparison of Reasoning-optimized and Chat LLMs on Ti-SafetyBench ($\times$100%)
LLM Version Method CA STEM Human Social Other China
DeepSeek V3 DA 38.3 38.52 39.7 39.61 38.2 35.46
CAA 37.9 37.93 38.89 39.45 38.2 35.03
R1 DA 43.75 44.81 38.57 47.39 44.24 43.73
CAA 47.01 46.97 44.68 49.79 48.23 45.37
GPT 4O DA 30.93 29.07 30.62 32.59 31.93 30.44
CAA 30.44 28.86 30.04 31.85 31.51 29.96
O1-mini DA 27.4 33.24 27.07 26.79 26.1 23.78
CAA 27.74 31.48 28.53 26.01 26.26 26.43
Table 12: Conditional Accuracy for Reasoning-optimized and Chat LLMs on Ti-MMLU ($\times$100%)
LLM Version Method Avg. STEM Human Social Other China
LlaMA 3.1-405B DA 25.08 23.88 24.25 25.58 27.62 24.07
CAA 25.28 24.10 24.50 25.87 27.73 24.22
3.1-70B DA 23.73 23.16 23.2 26.2 24.65 21.45
CAA 23.79 23.22 23.24 26.31 24.65 21.52
3.1-8B DA 5.47 5.48 5.56 5.99 5.46 4.86
CAA 7.44 7.95 7.54 7.38 7.41 6.92
Qwen 2.5-72B DA 7.27 6.07 7.98 7.52 7.74 7.02
CAA 16.50 15.73 17.88 17.00 15.84 16.04
2.5-32B DA 13.94 12.63 14.98 14.92 13.71 13.44
CAA 18.56 16.66 20.3 19.72 17.47 18.67
2.5-7B DA 1.8 2.94 1.87 1.63 0.9 1.68
CAA 14.59 13.92 13.66 16.34 14.57 14.46
Random - - 25.00 25.00 25.00 25.00 25.00 25.00
Table 13: Accuracy Comparison of LLMs Scales on Ti-MMLU ($\times$100%)
Model Version Method Avg. OFF UB PH MH IA EM PP
LlaMA 3.1-405B DA 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
CAA 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
3.1-70B DA 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
CAA 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
3.1-8B DA 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
Qwen 2.5-72B DA 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
CAA 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
2.5-32B DA 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
CAA 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
2.5-7B DA 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
CAA 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Random - - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 14: Accuracy Comparison of LLMs Scales on Ti-SafetyBench ($\times$100%)
Category GPT Claude Gemini DeepSeek LlaMA Qwen
3.5-turbo 4O O1-mini 3.5-Sonnet 1.5-flash v3 R1 3.1-8B 3.1-70B 3.1-405B 2.5-7B 2.5-32B 2.5-72B
business_ethics 3.35 22.01 10.53 42.11 38.76 33.97 30.62 6.22 23.44 24.40 15.79 14.83 10.05
economics 1.89 13.21 6.29 38.99 33.33 39.62 39.62 11.32 32.70 20.13 15.72 19.5 6.29
education 3.07 18.40 7.36 42.33 37.42 36.81 31.29 9.82 19.63 28.83 14.11 15.34 14.11
college_education 6.54 18.69 14.95 42.99 42.06 39.25 32.71 11.21 27.10 39.25 14.02 24.30 15.89
journalism 3.49 19.19 7.56 39.53 36.63 33.14 30.81 5.81 26.74 26.74 22.09 16.28 12.21
marketing 4.44 20.56 12.22 45.0 43.33 39.44 37.22 7.22 31.11 32.22 13.89 24.44 17.22
professional_accounting 5.14 20.00 9.14 41.14 37.71 33.71 30.86 6.86 20.57 22.86 18.86 19.43 12.00
professional_psychology 4.31 21.98 9.05 37.07 34.05 37.93 33.19 6.03 25.00 21.12 11.21 23.28 13.79
public_relations 2.87 11.49 10.34 32.18 33.33 36.78 29.89 5.17 24.14 24.14 17.82 20.69 13.79
security_study 5.19 25.19 10.37 42.96 37.78 38.52 25.93 7.41 30.37 26.67 20.00 20.00 14.07
high_school_geography 0.85 25.42 7.63 40.68 27.97 31.36 32.20 7.63 23.73 22.88 16.95 14.41 11.86
management 4.76 20.00 14.76 40.48 39.05 33.81 39.52 6.19 24.76 24.76 12.86 20.00 7.14
sociology 1.33 18.58 8.85 43.81 36.28 37.61 30.09 7.52 26.11 25.22 16.81 19.47 11.06
electrical_engineering 1.74 19.19 8.14 31.98 29.65 27.91 29.65 8.14 29.65 26.16 14.53 12.79 14.53
college_actuarial_science 4.72 6.60 8.49 18.87 17.92 25.47 24.53 8.49 17.92 16.04 5.66 17.92 12.26
college_engineering_hydrology 5.66 21.70 12.26 36.79 33.02 31.13 24.53 6.60 29.25 28.30 15.09 20.75 8.49
genetics 1.70 15.91 10.80 30.68 28.41 25.57 13.64 7.39 23.86 21.59 11.93 13.64 6.25
astronomy 3.03 11.52 16.36 35.76 23.03 30.91 24.24 7.27 20.00 26.06 13.94 16.36 11.52
college_mathematics 6.67 11.43 3.81 20.95 19.05 28.57 9.52 8.57 18.10 18.10 13.33 10.48 13.33
college_medical_statistics 5.66 14.15 15.09 41.51 34.91 28.30 5.66 11.32 24.53 27.36 16.98 25.47 9.43
virology 3.55 13.02 14.20 34.91 26.04 25.44 21.30 2.96 27.81 26.63 14.79 18.93 9.47
computer_science 4.41 14.71 11.27 25.49 29.41 29.41 20.10 8.33 24.51 28.92 10.78 15.69 9.80
conceptual_physics 0.68 16.33 10.88 41.50 27.89 25.85 30.61 13.61 23.13 27.89 14.29 15.65 13.61
anatomy 1.35 13.51 5.41 27.03 27.70 23.65 15.54 4.05 29.05 20.95 16.89 10.81 11.49
machine_learning 0.82 11.48 6.56 29.51 27.87 22.13 18.03 6.56 21.31 24.59 9.84 12.30 6.56
high_school_biology 5.33 17.16 7.69 20.12 24.26 22.49 8.28 9.47 18.34 20.71 8.88 12.43 8.28
high_school_chemistry 4.55 12.12 9.09 20.45 12.88 16.67 6.06 13.64 19.70 19.70 7.58 12.12 6.06
high_school_mathematics 6.71 12.20 9.15 34.15 25.00 37.20 27.44 7.32 20.12 24.39 26.83 24.39 18.90
high_school_physics 0.91 18.18 9.09 39.09 31.82 32.73 33.64 4.55 30.00 23.64 20.00 20.91 10.91
elementary_mathematics 7.39 13.04 6.52 36.09 34.78 26.09 44.35 6.96 17.39 28.70 15.22 22.61 14.78
legal_and_moral_basis 3.27 25.7 8.88 60.75 50.93 53.27 53.27 7.94 35.51 35.98 24.30 21.96 14.02
computer_security 3.51 22.81 12.87 44.44 32.16 30.99 25.15 8.77 30.99 30.41 16.37 17.54 9.94
food_science 2.10 20.28 6.29 34.27 25.17 37.76 25.17 5.59 28.67 32.87 14.69 18.88 11.19
college_medicine 1.47 16.12 6.59 26.37 21.25 23.81 12.45 8.06 22.71 19.05 10.26 12.09 9.16
clinical_knowledge 4.64 16.03 7.59 24.89 18.99 9.70 17.3 6.33 18.99 24.89 10.97 12.24 6.75
professional_medicine 3.46 13.83 10.64 23.67 18.88 24.47 19.15 10.90 17.82 18.09 11.44 16.76 8.24
human_sexuality 3.17 19.05 10.32 35.71 34.92 43.65 32.54 6.35 30.95 27.78 15.87 19.05 15.08
agronomy 2.96 15.38 8.28 31.95 30.18 33.14 28.99 5.33 15.98 26.04 12.43 18.34 10.65
sports_science 4.24 23.64 11.52 33.33 33.33 31.52 29.70 7.88 24.85 34.55 12.12 25.45 13.33
nutrition 2.07 11.72 13.79 37.24 37.24 31.72 20.69 6.90 20.00 27.59 17.24 12.41 7.59
elementary_information_and_technology 5.46 24.37 10.50 42.02 41.60 37.82 23.53 10.08 29.41 33.19 13.45 18.07 11.34
marxist_theory 1.06 21.16 10.05 48.68 38.62 41.27 40.21 10.05 32.28 23.81 16.40 14.81 16.93
college_law 3.70 18.52 9.26 26.85 18.52 19.44 32.41 10.19 22.22 22.22 11.11 17.59 12.04
global_facts 2.01 19.46 10.07 34.90 31.54 38.26 18.79 8.05 19.46 20.81 10.74 18.79 9.40
international_law 3.24 19.46 10.27 36.22 34.59 25.41 5.95 6.49 21.08 22.16 14.05 25.41 16.76
jurisprudence 2.92 19.71 9.49 39.17 30.41 37.71 32.36 9.25 21.65 27.25 15.09 19.22 11.44
world_religions 5.00 13.12 5.62 45.62 31.87 39.38 40.62 5.00 21.25 28.12 16.25 18.75 11.25
logical 3.25 15.45 10.57 38.21 28.46 40.65 26.83 5.69 26.02 30.08 12.20 23.58 12.20
professional_law 1.90 17.06 10.43 27.96 19.43 29.38 25.12 5.69 17.06 19.43 10.90 16.59 10.43
philosophy 3.81 19.05 13.33 44.76 38.10 42.86 1.90 7.62 28.57 27.62 14.29 20.00 10.48
world_history 4.97 15.53 11.80 36.02 28.57 30.43 32.30 7.45 24.84 24.22 16.15 22.36 16.15
arts 5.00 16.25 6.88 33.75 36.25 35.62 29.38 7.50 21.25 23.75 13.12 26.25 13.12
ancient_chinese 3.66 13.41 10.98 30.49 26.22 26.22 25.61 4.88 14.02 15.24 17.68 21.34 11.59
chinese_civil_service_exam 3.12 13.12 5.62 24.38 21.25 28.75 25.00 5.62 19.38 20.62 11.25 14.37 11.88
chinese_driving_rule 2.29 19.08 25.19 50.38 49.62 39.69 16.03 4.58 33.59 16.79 12.98 12.21 44.27
chinese_food_culture 2.94 19.12 6.62 27.94 34.56 29.41 36.76 4.41 29.41 24.26 11.76 17.65 15.44
chinese_foreign_policy 0.93 20.56 9.35 47.66 38.32 34.58 28.04 12.15 32.71 29.91 17.76 18.69 14.02
chinese_history 2.17 13.62 6.81 32.20 23.22 37.46 27.55 5.26 26.32 25.39 12.69 17.96 13.93
chinese_literature 2.45 18.63 13.73 23.04 28.43 29.41 26.47 7.35 19.61 21.57 15.20 19.61 14.71
chinese_teacher_qualification 2.23 17.32 10.06 45.81 36.87 25.14 41.34 6.15 22.91 34.08 15.64 20.11 13.41
construction_project_management 2.16 14.39 7.91 38.13 29.50 28.06 28.06 8.63 12.95 23.74 16.55 19.42 12.23
elementary_chinese 3.17 12.30 7.54 25.79 23.41 32.54 29.76 6.35 13.89 17.06 9.92 19.44 15.87
elementary_commonsense 6.06 16.16 8.08 39.90 32.83 35.86 36.36 6.06 24.24 28.28 14.14 20.20 9.60
ethnology 3.70 25.93 12.59 34.07 37.78 34.81 41.48 5.93 23.70 25.93 15.56 22.22 13.33
high_school_politics 4.20 20.28 6.99 30.07 23.08 30.77 26.57 6.29 21.68 26.57 15.38 18.18 10.49
modern_chinese 5.17 18.97 6.03 24.14 22.41 20.69 29.31 8.62 12.93 19.83 14.66 20.69 10.34
traditional_chinese_medicine 3.24 17.84 12.43 27.03 28.11 24.86 19.46 8.65 17.30 21.08 12.97 16.76 13.51
Table 15: Accuracy on 67 Subjects of Ti-MMLU (CAA) ($\times$100%)
Category GPT Claude Gemini DeepSeek LlaMA Qwen
3.5-turbo 4O O1-mini 3.5-Sonnet 1.5-flash v3 R1 3.1-8B 3.1-70B 3.1-405B 2.5-7B 2.5-32B 2.5-72B
business_ethics 1.91 20.57 4.78 40.67 36.84 30.14 19.14 5.74 23.44 24.4 1.44 12.44 7.66
economics 1.26 11.95 2.52 37.11 32.7 37.74 18.24 8.18 32.7 18.87 3.14 13.21 3.77
education 2.45 15.95 6.13 41.1 36.2 34.36 17.79 6.75 19.63 28.83 0.61 11.66 8.59
college_education 3.74 18.69 10.28 42.06 41.12 36.45 20.56 8.41 27.1 39.25 2.8 16.82 9.35
journalism 1.16 18.02 6.40 39.53 34.88 31.4 20.35 4.65 26.74 26.74 1.74 13.37 8.72
marketing 2.22 20.0 11.67 43.89 42.78 37.22 16.67 6.11 31.11 31.11 0.56 18.89 10.0
professional_accounting 4.57 17.14 6.29 39.43 37.71 30.86 15.43 6.29 20.57 22.29 0.57 13.71 5.14
professional_psychology 2.59 18.97 6.90 36.64 34.05 35.78 16.81 4.74 25.0 21.12 0.86 18.53 6.47
public_relations 1.15 9.77 8.62 32.18 32.18 33.91 18.39 5.17 23.56 24.14 1.72 17.24 5.17
security_study 3.7 22.22 5.93 40.0 37.04 34.07 12.59 5.19 29.63 26.67 3.7 14.07 14.81
high_school_geography 0.85 23.73 5.08 37.29 27.12 27.97 20.34 5.93 23.73 22.88 1.69 11.02 7.63
management 2.86 18.1 11.43 40.0 38.1 32.86 21.43 5.71 24.76 24.29 0.48 13.81 7.14
sociology 0.44 17.26 6.19 42.92 36.28 35.4 16.81 5.75 26.11 25.22 0.88 15.93 4.42
electrical_engineering 1.74 18.6 5.23 30.81 29.65 26.16 22.09 5.81 29.65 25.58 0.58 8.14 8.14
college_actuarial_science 2.83 4.72 6.60 16.98 16.98 22.64 24.53 7.55 17.92 16.04 3.77 13.21 8.49
college_engineering_hydrology 3.77 19.81 7.55 36.79 33.02 30.19 17.92 2.83 29.25 28.3 0.94 16.98 6.6
genetics 1.7 14.2 7.39 28.41 27.84 23.3 6.25 6.82 23.86 21.59 0.57 12.5 6.25
astronomy 2.42 10.3 9.7 34.55 21.82 26.06 16.97 6.67 20.0 26.06 0.61 13.94 5.45
college_mathematics 2.86 9.52 2.86 17.14 16.19 20.0 7.62 5.71 18.1 18.1 6.67 7.62 9.52
college_medical_statistics 2.83 14.15 4.72 41.51 33.02 22.64 2.83 3.77 24.53 26.42 0.94 22.64 3.77
virology 1.78 11.83 11.83 34.32 25.44 21.89 7.69 1.78 27.81 26.04 0.0 11.24 4.14
computer_science 2.45 12.75 4.41 25.0 28.43 27.45 11.27 4.9 24.51 28.43 0.49 11.27 4.41
conceptual_physics 0.68 14.97 7.48 35.37 25.85 22.45 12.24 10.2 23.13 27.89 1.36 11.56 3.4
anatomy 1.35 12.16 4.05 26.35 27.7 19.59 8.78 2.7 29.05 20.27 1.35 8.78 3.38
machine_learning 0.82 10.66 3.28 23.77 27.05 20.49 10.66 4.1 21.31 24.59 0.82 8.2 3.28
high_school_biology 4.14 15.38 6.51 11.24 24.26 18.34 1.78 7.1 18.34 20.71 1.18 7.69 4.14
high_school_chemistry 0.76 9.09 6.82 13.64 10.61 14.39 3.03 10.61 19.7 19.7 1.52 3.79 4.55
high_school_mathematics 5.49 11.59 6.71 31.71 25.0 33.54 23.78 5.49 20.12 24.39 17.68 22.56 13.41
high_school_physics 0.91 14.55 4.55 37.27 30.91 27.27 20.91 2.73 29.09 23.64 5.45 14.55 6.36
elementary_mathematics 4.35 12.17 4.78 31.74 33.91 24.35 37.39 4.35 17.39 28.26 6.09 20.0 7.83
legal_and_moral_basis 1.4 21.96 6.54 58.88 48.6 50.47 34.11 7.48 35.51 35.51 0.93 17.76 10.28
computer_security 1.75 20.47 8.77 39.18 30.41 29.24 12.28 7.6 30.99 30.41 3.51 14.04 8.77
food_science 2.10 19.58 1.40 32.87 23.78 32.17 13.29 2.8 28.67 32.87 0.00 16.08 6.99
college_medicine 0.73 15.38 3.66 26.01 21.25 22.71 6.23 5.49 22.71 19.05 1.47 6.96 3.66
clinical_knowledge 3.38 15.61 5.91 24.89 17.72 9.28 5.49 4.22 18.99 24.89 1.27 10.55 8.44
professional_medicine 2.66 12.23 7.18 22.87 18.62 23.14 6.91 8.24 17.82 18.09 0.53 12.5 6.91
human_sexuality 2.38 17.46 6.35 34.13 34.92 42.06 16.67 3.97 30.95 27.78 0.0 16.67 10.32
agronomy 0.59 13.02 5.33 31.36 30.18 30.18 20.71 4.73 15.98 25.44 0.0 13.02 7.69
sports_science 2.42 22.42 7.88 33.33 33.33 30.30 13.33 7.27 24.85 34.55 0.61 21.21 10.91
nutrition 1.38 11.03 7.59 36.55 37.24 29.66 10.34 2.76 20.0 27.59 0.69 8.28 3.45
elementary_information_and_technology 3.78 21.01 7.98 39.92 41.18 35.71 13.87 5.88 29.41 33.19 2.1 14.29 4.20
marxist_theory 0.53 19.58 7.41 48.15 37.57 39.68 22.22 6.88 31.75 23.81 2.65 10.05 8.47
college_law 1.85 15.74 8.33 25.93 18.52 16.67 17.59 6.48 22.22 21.3 1.85 12.04 6.48
global_facts 0.67 18.12 6.71 33.56 30.87 38.26 14.09 7.38 19.46 20.81 1.34 16.11 8.05
international_law 2.16 19.46 5.95 35.14 34.05 24.32 2.70 5.41 21.08 22.16 1.62 21.08 9.73
jurisprudence 1.95 17.52 7.06 35.77 27.98 36.25 17.52 8.03 21.65 27.01 1.95 12.41 9.25
world_religions 2.5 11.88 3.12 45.62 31.87 35.0 12.2 2.5 21.25 27.5 0.62 13.12 7.5
logical 2.44 15.45 8.13 36.59 27.64 39.02 12.20 4.07 26.02 30.08 2.44 17.07 5.69
professional_law 0.95 16.11 4.74 26.07 18.96 25.12 7.58 4.74 17.06 18.48 0.0 9.95 5.69
philosophy 2.86 18.1 5.71 44.76 37.14 40.0 0.95 5.71 28.57 27.62 1.9 16.19 7.62
world_history 3.73 13.04 8.07 34.78 26.09 27.33 16.77 4.97 24.84 24.22 3.11 17.39 10.56
arts 4.38 15.0 4.38 33.12 36.25 30.00 17.50 5.00 21.25 23.75 3.12 19.38 8.75
ancient_chinese 1.83 12.2 5.49 28.05 26.22 22.56 11.59 3.05 13.41 15.24 1.83 15.85 5.49
chinese_civil_service_exam 0.62 10.0 2.5 19.38 18.12 25.0 15.62 3.75 19.38 20.62 3.75 8.75 7.5
chinese_driving_rule 0.76 16.79 7.63 50.38 49.62 38.93 19.85 3.82 33.59 24.43 0.76 11.45 6.11
chinese_food_culture 2.21 17.65 2.94 27.21 33.82 27.94 18.38 4.41 29.41 24.26 0.0 13.97 7.35
chinese_foreign_policy 0.0 20.56 6.54 42.99 36.45 32.71 18.69 10.28 32.71 28.97 2.8 12.15 7.48
chinese_history 1.55 11.46 3.72 30.65 21.98 31.58 12.38 4.64 26.32 25.08 1.55 13.62 6.19
chinese_literature 1.47 17.65 6.86 22.55 28.43 27.45 19.61 4.9 19.61 21.08 1.96 16.18 9.31
chinese_teacher_qualification 2.23 16.76 5.03 44.13 35.75 24.02 24.58 4.47 22.91 34.08 1.12 15.64 8.94
construction_project_management 1.44 12.95 5.04 35.97 28.06 25.9 17.27 5.04 12.95 23.74 1.44 10.07 5.76
elementary_chinese 2.38 11.9 3.97 24.21 23.41 29.37 18.65 4.76 13.49 17.06 1.19 13.89 8.33
elementary_commonsense 3.54 15.66 3.54 39.9 32.83 32.83 24.24 3.54 24.24 28.28 2.02 15.66 5.56
ethnology 2.22 22.96 5.19 32.59 36.3 30.37 33.33 4.44 23.7 25.93 2.22 18.52 12.59
high_school_politics 2.1 18.18 2.10 26.57 18.88 28.67 20.98 5.59 21.68 26.57 2.8 14.69 3.5
modern_chinese 3.45 18.97 6.03 20.69 21.55 18.97 14.66 3.45 12.93 19.83 1.72 12.07 2.59
traditional_chinese_medicine 1.62 15.68 7.03 27.03 27.57 24.32 9.73 4.86 17.3 21.08 1.08 10.81 7.03
Table 16: Accuracy on 67 Subjects of Ti-MMLU (DA) ($\times$100%)
Appendix D Appendix: Qualitative Analysis
To complement our quantitative results, we provide a brief qualitative analysis of representative model failures in TLUE tasks. As shown in Table 10, the errors primarily fall into two categories: knowledge-related mistakes and repetitive output behaviors. For instance, Claude-3.5-Sonnet Anthropic (2024) incorrectly assessed the physical properties of “colorless film,” while GPT-3.5-Turbo Achiam et al. (2023) simply repeated the input without reasoning. In contrast, GPT-4O Achiam et al. (2023) correctly handled the same inputs, demonstrating more robust comprehension. These cases highlight common pitfalls in model understanding under low-resource settings.
Figure 10: Bad Case Analysis
As shown in Figure. 11, interestingly, Claude-3.5-Sonnet Anthropic (2024) experienced a significant performance drop under the 5-shot setting, as it erroneously reproduced the answers from all five in-context examples along with the test answer—a unique behavior not observed in other models. We believe this reflects issues in instruction-following for Tibetan, and we provide representative examples of this failure case in the main text.
Figure 11: Bad Case Analysis
To identify the root causes of model failures on TLUE, we conducted a qualitative analysis under consistent zero-shot prompting. Errors were mainly categorized as: (1) Knowledge Errors, and (2) Repeater Behaviors. These patterns suggest that failures are due to model limitations, rather than prompt misinterpretation.
Figure 12: Bad Case Analysis
Appendix E Appendix: Few-Shot Evaluation
We have further explored the few-shot setting, with the results presented in Table 17. These findings demonstrate the effectiveness of few-shot prompting and underscore the potential of adapting language models for Tibetan.
Notably, LLMs that performed below the random baseline in the zero-shot setting showed marked improvements, suggesting their initial under performance may be attributed to task misinterpretation. In contrast, stronger models such as Gemini-1.5-Flash DeepMind (2024) and DeepSeek-V3 DeepSeek-AI et al. (2024) exhibited only marginal gains, indicating their challenges lie in genuine linguistic understanding rather than prompt misalignment.
LLM Version Average Score (0shot/5shot) Improvement
Claude 3.5-Sonnet 35.63 / 13.57 ↓22.06
Gemini 1.5-Flash 31.01 / 32.67 ↑1.66
DeepSeek v3 32.16 / 33.63 ↑1.47
GPT 4O 17.51 / 23.56 ↑6.05
3.5-Turbo 3.42 / 18.51 ↑15.09
LlaMA 3.1-70B 23.79 / 26.92 ↑3.13
3.1-8B 7.44 / 20.09 ↑12.65
Qwen 2.5-7B 14.59 / 22.79 ↑8.20
2.5-32B 18.56 / 23.78 ↑5.22
2.5-72B 16.50 / 27.99 ↑11.49
Table 17: LLM Performance Comparison (0 Shot/ 5 shot) with Improvement ($\times$100%)
Appendix F Appendix: Statistical Significance and Robustness Analysis
We also conducted 10 independent runs for each model and computed average scores along with their standard deviations and 95% confidence intervals.
The Table 18 summarizes these statistics for several models evaluated on the Ti-MMLU benchmark. Including error bars and confidence intervals offers a clearer understanding of the variability and reliability of the reported accuracy metrics.
LLM Version AVG. (CAA) Std. Dev. Lower 95% CI Upper 95% CI
Claude 3.5-Sonnet 35.63 1.20 33.28 37.98
Gemini 1.5-Flash 31.01 1.00 29.05 32.97
GPT O1-mini 9.67 0.50 8.70 10.64
4O 17.51 0.80 16.05 18.97
3.5-Turbo 3.42 0.30 2.85 3.99
DeepSeek R1 27.45 1.05 25.40 29.50
V3 32.16 1.10 29.99 34.33
LlaMA 3.1-405B 25.28 0.90 23.52 27.04
3.1-70B 23.79 0.85 22.13 25.45
3.1-8B 7.44 0.40 6.67 8.21
Qwen 2.5-32B 18.56 0.70 17.10 20.02
2.5-72B 16.50 0.65 15.23 17.77
2.5-7B 14.59 0.60 13.42 15.76
Table 18: LLM Performance with Confidence Interval (CI) ($\times$100%)
Appendix G Appendix: DA and CAA Evaluation Metrics
To evaluate model performance in low-resource Tibetan settings, we use two complementary metrics: DA and CAA.
DA extracts the first uniquely occurring A/B/C/D option from the model’s output, reflecting the model’s ability to follow Tibetan instructions precisely. CAA, in contrast, handles cases where LLMs list all options or include reasoning. It filters such outputs and identifies the correct answer, offering an upper-bound estimate. The DA Result is shown in Figure. 13, and the CAA result is shown in Figure. 14.
Figure 13: Bad Case Analysis
Figure 14: Bad Case Analysis
Appendix H Appendix: Clarification on Model Comparison under Low Accuracy
While some LLMs score below the random baseline in overall accuracy, our evaluation still provides informative signals at the instance level. This is especially relevant in zero-shot settings for Tibetan, where correctness alone cannot fully capture LLM behavior.
Each prediction reflects whether a LLM can handle specific domains or linguistic constructions. To complement the quantitative findings in Section 5.6, we include a qualitative comparison in Table 15. This example contrasts two LLMs from the Qwen-2.5 series Team (2024) on the same Ti-MMLU question. Notably, the larger 72B model generates a fluent but incorrect explanation, while the 32B model selects the correct answer with a more concise yet accurate rationale. This case illustrates non-monotonic scaling behavior and differences in reasoning stability, highlighting the value of instance-level analysis even under low average performance.
Figure 15: Prompt Sensitivity Analysis
Appendix I Appendix: Prompt Sensitivity Analysis
We also evaluated multiple prompt variants to assess whether minor differences in phrasing impact LLM performance. The results indicate that prompt wording has a negligible effect on LLM predictions. An illustrative example across four prompt templates is provided in Figure. 16.
Figure 16: Prompt sensitivity Analysis